### v1

# Assessing the dynamical assumptions in Tsirelson inequality tests of non-classicality in harmonic oscillators

Abstract
“Macrorealism” posits that a system possesses definite properties at all times and that we can discover these properties, in principle, without disturbing the system’s subsequent behaviour. The Leggett–Garg inequalities are derived under these assumptions and are readily violated by standard quantum mechanics, thereby providing a scheme to test whether demonstrably macroscopic systems can exhibit quantum coherence. Unfortunately, Leggett–Garg tests suffer from the ineludible clumsiness loophole—the difficulty of proving that sequential measurements have not inadvertently disturbed the system. The recently uncovered Tsirelson inequality is derived from the simple dynamical assumption of uniform precession, obeyed by many classical systems, and requires only single-time measurements. However, Tsirelson inequality violations could be explained by a macrorealistic system that merely breaks the dynamical assumption, rather than genuine quantum behaviour. By carrying out a quantum-mechanical analysis of the Tsirelson inequality in the harmonic oscillator, we develop a protocol to rule out this possibility by assessing generalised conditions of uniform precession. We show that various measures of uniform precession, some of which related to Leggett–Garg quantities, are satisfied well enough that the presence of quantum-mechanical interference terms must be implied. We derive several incidental mathematical results relating to maximally-violating states, concerning dwell time, crossing number and probability currents, and also consider a group theoretic analysis of the Tsirelson operator.
preprint: APS/123-QED
I Introduction
I.1 The Tsirelson inequality
In 2006, Boris Tsirelson wrote a paper with a title that asks an endearingly straightforward question: “How often is the coordinate of a harmonic oscillator positive?” [1]. Tsirelson demonstrated that a simple dynamical assumption about classical oscillators can give rise to a surprising disparity between classical and quantum expectations. Violations of the “Tsirelson inequality” [2] are associated inherently with the negativity of the Wigner function [3], and present an alluring scheme to assess the “quantumness” of a system using only single-time measurements on an ensemble of systems.
This work went virtually unnoticed for more than a decade and a half, until it was rediscovered by Zaw et al., who in their 2022 paper extended Tsirelson’s work to a broader family of systems undergoing uniform precession, such as finite-dimensional spins in real space [4]. Tsirelson’s protocol has been generalised as a viable entanglement witness for coupled harmonic oscillator systems [5, 6] and entangled spin ensembles [7], and the most recent work in the field has focussed primarily upon further generalisations of the protocol [8] and computing stricter bounds on the maximal quantum violations of Tsirelson’s inequality [9, 10]. A recent triumph has been the first experimental realisation (a so-called quantumness certification) of Tsirelson violations in precessing nuclear spins [11].
To derive the Tsirelson inequality, we consider a classical harmonic oscillator with time period $T$. We define the dichotomous variable $Q(t)$ to be the sign function $\text{sign}(x(t))$, which is $\pm 1$ depending on the sign of $x$, and measure this function—which is the mathematical equivalent of asking the question, on which side of the origin does the particle lie?—at three equally-spaced times, every one-third of a period ($t_{1}=0$, $t_{2}=\frac{1}{3}T$, and $t_{3}=\frac{2}{3}T$). A simple dynamical assumption about oscillators is this: the particle will never be found on the same side of the origin three times in a row. Therefore, the sum $Q(0)+Q(\frac{1}{3}T)+Q(\frac{2}{3}T)$ can only assume the values +1+1 or 1-1 and not +3+3, for instance, because this would require $Q=+1$ at all three times. This is illustrated in Fig. 1. (Tsirelson originally considered the harmonic oscillator’s classical phase-space dynamics, illustrated in Fig. 2, and the probability that a particle is found on one side of the origin.)
Figure 1: A classical particle oscillating in a quadratic potential is one example of a system undergoing uniform precession. The dynamical assumption is that the value of $Q$ must take the opposite sign at one of the three measurement times, irrespective of where in its trajectory the particle starts at time $t=0$.
Defining $A=Q(0)+Q(\frac{1}{3}T)+Q(\frac{2}{3}T)$, it follows that the ensemble average $\langle A\rangle$ over several experimental runs must satisfy the Tsirelson inequality:
$1\leq\langle A\rangle\leq 1\ .$ (1.1)
Since the aforementioned dynamical assumption holds true for any system undergoing uniform precession, this Tsirelson inequality would be respected by even, for example, a (classical) charged particle precessing in a magnetic field. However, we focus on the case of the harmonic oscillator for the remainder of this work.
By virtue of the classical equivalence between the expectation value of a sum and the sum of expectation values, $\langle A\rangle$ can just as well be written in terms of single-time averages:
$\langle A\rangle=\langle Q(0)\rangle+\langle Q(\frac{1}{3}T)\rangle+\langle Q(\frac{2}{3}T)\rangle$ (1.2)
Thus, the Tsirelson quantity $\langle A\rangle$ can be measured using only single-time measurements on an ensemble of identical systems. In practice, measuring the dichotomic variable $Q$, without ascertaining any further information about the particle’s position, may be done using a light beam that illuminates one half of the harmonic well [12].
However, a crucial point which has remained hitherto unaddressed in the literature is the validity of the dynamical assumption of uniform precession in a quantum-mechanical context. If truly the Tsirelson test is to be interpreted as a credible certification of quantum behaviour, it is imperative to demonstrate that violations are indicative of genuine quantum phenomena and not of simply the failure of its founding dynamical assumption.
Figure 2: Classical phase-space depiction of the system’s dynamics for three possible measurement times. If the system is initialised in the green region for one of the experimental runs, then it will have a probability of $2/3$ of being found on the right side of the origin at a randomly-chosen measurement time $t\in{0,T/3,2T/3}$. Conversely, the probability will be $1/3$ if the particle begins in a blue region. As such, the ensemble average, a convex sum of these probabilities, is bounded between $\frac{1}{3}$ and $\frac{2}{3}$. This is equivalent to Eq. 1.1 up to a scale factor. (Adapted from Zaw et al. [13].)
Uniform precession can be defined perfectly well in the case of a classical oscillator by considering, for instance, the phase-space representation of its trajectory. The condition for uniform precession is then simply that the angular variable $\varphi=\tan^{-1}(x/p)$ evolve linearly in time ($\dot{\varphi}=\text{const.}$). However, quantum-mechanically, the corresponding operator, related to the arrival time operator, is conjugate to the Hamiltonian and is not self-adjoint [14, 15, 16, 17]. As such, the eigenstates of the angular operator cannot be orthogonal, and there is some intrinsic indefiniteness in the angular behaviour and hence in the definition of uniform precession. This property is also related to the difficulties of defining a phase operator for the harmonic oscillator [18]. In practice, this means that uniform precession can be defined in multiple inequivalent ways in quantum mechanics, as we shall see.
The Tsirelson quantity $\langle A\rangle$ may be conveniently written in the Wigner representation [3, 19, 20, 21, 22, 23, 9] in terms of the function $W(p,x,t)$ (and see also Appendix A). It has the form
$\langle A\rangle=\iint\text{sign}(x)\sum_{i}W(p,x,t_{i})\ \text{d}x\ \text{d}p\ .$ (1.3)
Although the Wigner function evolves along the classical path, it is not a probability distribution, since it can go negative, and thus we cannot assert that the evolution in the quantum case respects uniform precession. In the Heisenberg picture with a time independent Wigner function, the Tsirelson quantity is
$\langle A\rangle=\iint\left[Q(0)+Q(\frac{T}{3})+Q(\frac{2T}{3})\right]W(p,x)\ \text{d}x\ \text{d}p\ ,$ (1.4)
which makes it clear that Tsirelson inequality violations are a negative Wigner function effect.
We dedicate the remainder of this paper to finding precise sets of conditions for uniform precession, in terms of measurable probabilities, that are all satisfied classically, but which lead to quantum-mechanically inequivalent expressions and measures, which we may use to assess the degree to which the dynamical assumption is satisfied. In order to do so, we will embed the Tsirelson scheme within the more comprehensive Leggett–Garg framework [24]—the canonical test of non-classicality in temporal correlations—and utilise some of its powerful machinery.
I.2 Leggett–Garg tests of non-classicality
The Leggett–Garg inequalities (LGIs) [24, 25], a temporal analogue to Bell-type inequalities [26, 27], must be satisfied by any theory obeying “macrorealism”, a quintessentially classical world-view that is encapsulated by the following three postulates: macroscopic realism per se—a macroscopic system exists in a definite state at all times; non-invasive measurability—the state of the system can be determined, in principle, without affecting its subsequent dynamics; and induction—the state does not depend upon future measurements.
Macrorealism pertains to measurements made on a specific data set, which we take here to be the set of three single-time averages $\langle Q_{i}\rangle$, where $Q_{i}\equiv Q(t_{i})$, and the three correlators $C_{ij}$, the latter measured in a non-invasive way. With possible measurement outcomes $s_{i}=\pm 1$, where $i=1,2,3$ for measurements taken at times $t_{1},t_{2},t_{3}$, respectively, the two-time correlators are given by
$C_{ij}\equiv\langle Q_{i}Q_{j}\rangle=\sum_{s_{i},s_{j}}s_{i}s_{j}p_{ij}(s_{i},s_{j})\ ,$ (1.5)
with $1\leq C_{ij}\leq 1$, and where $p_{ij}(s_{i},s_{j})$ is the proportion of outcomes in the data set where a measurement of $Q$ at time $t_{i}$ yields $s_{i}$ and at (a later time) $t_{j}$ yields $s_{j}$.
It will be useful to introduce here some mathematical expressions we will require for later analysis. The measured data set fully determines the two-time quasi-probability distribution [28, 29]
$q(s_{i},s_{j})=\frac{1}{4}\left(1+s_{i}\langle Q_{i}\rangle+s_{j}\langle Q_{j}\rangle+s_{i}s_{j}C_{ij}\right)\ ,$ (1.6)
for $s_{i},s_{j}=\pm 1\ ,\ j>i\in\{1,2,3\}$, and partially determines the three-time quasi-probability
$q(s_{1},s_{2},s_{3})=\frac{1}{8}(1+s_{1}\langle Q_{1}\rangle+s_{2}\langle Q_{2}\rangle+s_{3}\langle Q_{3}\rangle\\ +s_{1}s_{2}C_{12}+s_{2}s_{3}C_{23}+s_{1}s_{3}C_{13}+s_{1}s_{2}s_{3}D)\ ,$ (1.7)
where $D$ is the triple correlator, $\langle Q_{1}Q_{2}Q_{3}\rangle$, but is never measured in the simplest protocols. These two quasi-probabilities are classically positive. Quantum-mechanically, they are equivalent to the expressions
$q(s_{1},s_{2})=\text{Re}\left\{\text{Tr}\left(P_{s_{2}}(t_{2})P_{s_{1}}(t_{1})\rho\right)\right\}\ ,$ (1.8)
$q(s_{1},s_{2},s_{3})=\text{Re}\left\{\text{Tr}\left(P_{s_{3}}(t_{3})P_{s_{2}}(t_{2})P_{s_{1}}(t_{1})\rho\right)\right\}\ ,$ (1.9)
with projection operators $P_{s_{i}}(t_{i})=\frac{1}{2}(1+s_{i}Q_{i})$. We also give here the analogous sequential measurement probabilities and their moment expansions:
$p_{12}(s_{1},s_{2})=\text{Tr}(P_{s_{2}}(t_{2})P_{s_{1}}(t_{1})\rho P_{s_{1}}(t_{1}))=\frac{1}{4}\left(1+s_{1}\langle Q_{1}\rangle+s_{2}\langle Q_{2}^{(1)}\rangle+s_{1}s_{2}C_{12}\right)\ ,$ (1.10)
$p_{123}(s_{1},s_{2},s_{3})=\text{Tr}(P_{s_{3}}(t_{3})P_{s_{2}}(t_{2})P_{s_{1}}(t_{1})\rho P_{s_{1}}(t_{1})P_{s_{2}}(t_{2}))\\ =\frac{1}{8}\left(1+s_{1}\langle Q_{1}\rangle+s_{2}\langle Q_{2}^{(1)}\rangle+s_{3}\langle Q_{3}^{(12)}\rangle+s_{1}s_{2}C_{12}+s_{2}s_{3}C_{23}^{(1)}+s_{1}s_{3}C_{13}^{(2)}+s_{1}s_{2}s_{3}D\right)\ ,$ (1.11)
where the superscripts denote the presence of an earlier or intermediate measurement that has been summed out. The difference between, for example, $\langle Q_{2}\rangle$ and $\langle Q_{2}^{(1)}\rangle$ signifies the presence of the interference terms which produce the LGI violations.
Turning now to the conditions for macrorealism, the three postulates of macrorealism imply that there exists an underlying joint probability for the given data set, which implies that a set of twelve two-time and four three-time LG inequalities hold. Specifically, the twelve quasi-probabilities Eq. 1.6—for time pairs $(t_{1},t_{2})$, $(t_{1},t_{2})$, $(t_{1},t_{2})$—are non-negative (LG2s) and the following four LG3 inequalities are satisfied:
$L_{1}=\frac{1}{4}(1+C_{12}+C_{23}+C_{13})\geq 0$ (1.12)
$L_{2}=\frac{1}{4}(1-C_{12}-C_{23}+C_{13})\geq 0$ (1.13)
$L_{3}=\frac{1}{4}(1-C_{12}+C_{23}-C_{13})\geq 0$ (1.14)
$L_{4}=\frac{1}{4}(1+C_{12}-C_{23}-C_{13})\geq 0\ .$ (1.15)
These sixteen conditions are not only necessary for macrorealism but also sufficient, as the LG version of Fine’s theorem shows [30, 31, 32]. In quantum mechanics, the LG2 and LG3 inequalities can be violated up to the value $\frac{1}{8}$, the Lüders bound. [^1] This is for so-called Lüders measurements, which simply measure the value of $Q$. More negative values can be obtained by measuring finer-grained quantities and then coarse-graining the probabilities to find probabilities of $Q$ [33, 34].
We also note that the four LG3s may be written as
$L_{1}=q(+,+,+)+q(-,-,-)$ (1.16)
$L_{2}=q(+,-,+)+q(-,+,-)$ (1.17)
$L_{3}=q(+,-,-)+q(-,+,+)$ (1.18)
$L_{4}=q(+,+,-)+q(-,-,+)\ ,$ (1.19)
from which we see that the LG3s describe a quasi-probability distribution for changes of sign.
This means that if the LG3s are satisfied but the LG2s are not, then even though MR is violated for the full data set, it can still be satisfied for the data subset consisting of the sign-change variables, $\chi_{1}=s_{1}s_{2}$ and $\chi_{2}=s_{2}s_{3}$, which are $\pm 1$ depending on whether or not the sign of $Q$ has flipped between measurement times. This is a point we will make extensive use of in Section III, where we are interested in the sign-change probabilities as a quantification of uniform precession.
Violations of the LGIs have been observed in a myriad of physical systems, such as superconducting qubits [35, 36], nuclear spins [37, 38, 39], millimetre-scale crystals [40], photons [41], and (recently) neutron interferometry [42, 43]. The Leggett–Garg inequalities also have far-reaching applications to other fields [44, 45, 46, 47, 48]. A more exhaustive account of the experimental tests of macrorealism to date is provided in the extensive review articles Refs. [33, 49].
However, a significant issue is that macrorealism per se is conjoined inextricably with non-invasive measurability in the Leggett–Garg scheme. The “clumsiness loophole” [50] is that any recorded violations—even if a system truly does not admit realism—could always be ascribed, instead, to the unwitting invasiveness of the measurement process. Indeed, a perfectly classical stochastic model with measurements that disturb the system can reproduce correlation functions that violate the LGIs [51]. The standard way to render the measurements non-invasive is to use ideal negative result measurements [52, 53] (and its refinements [54]), but alternative methods exist, such as the waiting detector protocol [55, 56, 57]. If the non-invasive measurement condition is implemented convincingly, then violations of the LGIs imply the failure of macrorealism per se.
The derivation of the Leggett–Garg inequalities solely from the postulates of macrorealism is completely independent of the system’s equations of motion. It is this indifference to the underlying dynamics that necessitates sequential measurements in order to elucidate a quantum–classical disparity. But the prospect of observing non-classicality in temporal correlations need not be plagued by the experimentally challenging task of circumventing invasive sequential measurements. The introduction of a judiciously chosen and experimentally testable dynamical assumption can supplant the requirement for non-invasive measurement. A few examples of such assumptions have appeared in the LG literature and typically consist of assuming that certain transition probabilities are zero [58, 59, 55] or can be simplified in some way, for example via a stationarity assumption [33, 60, 61]. The key point is that the assumptions are chosen so that the correlators simplify to single-time measurement averages. Furthermore, the dynamical assumption is either implemented through the design of the experiment (e.g. the choice of Hamiltonian, where possible) or checked experimentally through separate control experiments. Dynamical assumptions are often classically motivated, and the Tsirelson scheme is an example par excellence, with its simplifying dynamical assumption of uniform precession.
The related single crossing dynamical assumption, studied in the context of a waiting detector model in Ref. [55] and Chapter 6 of Ref. [57], is mathematically equivalent over half a period to the assumption of uniform precession over a full period, an elegant proof of which we provide in Appendix B.
But does a violation truly imply genuine quantum behaviour and the failure of macrorealism, or could a stubborn macrorealist now argue that it stems purely from faulty dynamics? A classical system that breaks the dynamical assumption—that is, one which is not precessing uniformly (or not oscillating at all)—will easily violate the Tsirelson inequality. Fig. 3 depicts qualitatively that a Tsirelson violation could be caused by a classical particle that does not obey the expected Newtonian dynamics of oscillation induced by a harmonic potential, lingering too long on one side of the origin; or perhaps we have even simply defined the origin inaccurately. Such a trajectory still has a defined position at all times, and hence is macrorealistic, yet a spurious Tsirelson inequality violation will ensue. So, just as separating realism from non–invasive measurability has proved to be the greatest hassle in LG tests, do we now run into a similar problem in Tsirelson tests of separating the failure of realism from this newfound dynamical assumption?
Figure 3: Unlike the blue classical trajectory, which follows the expected sinusoidal dynamics, the erractic red trajectory, still macrorealistic, dwells for too long on the right side of the origin and will violate the Tsirelson inequality, despite there being no quantum behaviour. This is the faulty dynamics loophole.
Our strategy to assess the dynamical assumption will be to re-write the Tsirelson quantity in terms of various probabilities for two and three time measurements that can be interpreted as measures of uniform precession. We will do this in two ways. Our first method involves demonstrating an explicit connection between $\langle A\rangle$ and the probability of zero sign changes, $L_{1}$ (Eqs. 1.12 and 1.16), which should be zero classically. Our second method stems from the interesting observation that a weaker condition of setting merely the difference q(+,+,+)q(,,)q(+,+,+)-q(-,-,-) to be zero also leads to the Tsirelson inequality, as we shall show, and thus may also be considered a measure of uniform precession.
I.3 This paper
Assessing how well the dynamical assumption of uniform precession holds and classifying the nature of Tsirelson violations hereinafter forms the core narrative of this paper. In Section II, we lay forth the mathematical machinery required for quantum-mechanical calculations and discuss quantum states and their violations of the Tsirelson inequality, proving some results relating to maximally-violating states through a group theoretic perspective. We present our most important results in Section III on the main protocol we have devised to assess uniform precession, broken down into the two methods of defining uniform precession discussed prior. Here we shall require the breadth of the Leggett–Garg formalism to disentangle violations of macrorealism from the breaking of uniform precession. In Section IV, we investigate two other quantities that are less strictly related to the Tsirelson inequality but provide further heuristic evidence of the extent to which uniform precession is satisfied. We will show that one has quantum properties identical to the classical case and the other approximately so. Moreover, we will argue that these results can be used to rule out alternative classical explanations. Finally, we conclude in Section V with a critical assessment of whether the protocol introduced in Section III ultimately sidesteps the problem of non-invasive measurement. We report lengthy technical results and rigorous mathematical proofs in the appendices.
II Tsirelson Violations in the Quantum Harmonic Oscillator
II.1 Expansion in energy eigenbasis
Tsirelson showed the unexpected result that probing a quantum harmonic oscillator (QHO) at different times can produce ostensibly non-classical results, despite the fact that the time evolution of quantum operators in the Heisenberg picture abides by the same classical equations of motion.
In order to phrase the problem quantum-mechanically, we promote the variable $Q$ to be an operator $\hat{Q}$ in the Heisenberg picture, undergoing unitary time evolution as $\hat{Q}(t)=U^{\dagger}\hat{Q}U$, where $U=\text{exp}(-itH)$ with a standard QHO Hamiltonian $H$. We can write $\hat{Q}=(2\hat{\theta}-1)$, where $\hat{\theta}$ is the Heaviside operator. For convenience, we will work exclusively with units chosen such that $\hbar=1$; further, it will be convenient to normalise the angular frequency of oscillations to unity, $\omega=1$. The Tsirelson inequality is then $1\leq\langle A\rangle\leq 1$, where now
$A\equiv\hat{Q}+U^{\dagger}\hat{Q}U+(U^{\dagger})^{2}\hat{Q}U^{2}\ ,$ (2.1)
with $U=\text{exp}(-\frac{2\pi}{3}iH)$. We will refer to $A$ hereafter as the “Tsirelson operator”. Quantum systems can violate this inequality—up to a maximal value of $\langle A\rangle\approx 1.26$, so Tsirelson found [1].
We can compute the Tsirelson quantity for an arbitrary state of the QHO Hilbert space formed as a superposition of energy eigenstates, from the ground state upwards, up to a finite cut-off:
$|\Psi\rangle=\sum_{n=0}^{N}c_{n}|n\rangle,c_{n}\in\mathbb{C}\ .$ (2.2)
The single-time average $\langle\hat{Q}\rangle=\langle\Psi|2\hat{\theta}-1|\Psi\rangle$ can be written as
$\langle\Psi|2\hat{\theta}-1|\Psi\rangle=\sum_{n,k=0}^{N}\langle\Psi|n\rangle\langle n|2\hat{\theta}-1|k\rangle\langle k|\Psi\rangle\\ =\sum_{n,k=0}^{N}c_{n}^{*}Q_{nk}c_{k}\ ,$ (2.3)
where we have defined the matrix elements $Q_{nk}=\langle n|2\hat{\theta}-1|k\rangle$. A tractable expression for $Q_{nk}$ is derived in Appendix C and is given by
$\frac{2\pi^{-\frac{1}{2}}}{(k-n)\sqrt{2^{n+k}n!k!}}[kH_{k-1}(0)H_{n}(0)-nH_{n-1}(0)H_{k}(0)]\ ,$ (2.4)
where $H_{n}(x)$ are the standard Hermite polynomials.
In the Schrödinger picture,
$|\Psi,t\rangle=\sum_{n=0}^{N}c_{n}e^{-in\omega t}|n\rangle\ ,$ (2.5)
since $|n\rangle$ is by definition an eigenstate of the QHO Hamiltonian with eigenvalue $\omega(n+\frac{1}{2})$. We can drop the vacuum contribution of $\frac{1}{2}\omega$ since this will have no effect upon the expectation value:
$\langle\hat{Q}(t)\rangle=\langle\Psi,t|\hat{Q}|\Psi,t\rangle=\sum_{n,k=0}^{N}c_{k}^{*}c_{n}e^{-i(n-k)\omega t}Q_{nk}\ .$ (2.6)
Therefore, for an arbitrary state up to an energy truncation of $N$, the Tsirelson quantity is given by
$\langle A\rangle=\sum_{n,k=0}^{N}c_{k}^{*}c_{n}(1+e^{-i\frac{2\pi}{3}(n-k)}+e^{-i\frac{4\pi}{3}(n-k)})Q_{nk}\ ,$ (2.7)
recalling that, by choosing units of $\omega=1$, one full period corresponds to $T=2\pi$.
Violations of the Tsirelson inequality are manifestly associated with negativity in the Wigner function. A corollary is that a single coherent state, which has a completely positive Wigner function, will not violate the Tsirelson inequality. Neither will a single energy eigenstate, for it is trivially of strictly odd or even parity; a neat proof, in terms of probability currents, that definite-parity states do not violate the Tsirelson inequality is provided in Section IV.1. Perhaps more surprising is that a superposition of at least three energy eigenstates is required to produce a violation. We provide a full proof of this fact in Appendix E.
II.2 Maximally violating eigenstates
The Tsirelson operator is Hermitian ($\hat{A}\equiv\hat{Q}_{1}+\hat{Q}_{2}+\hat{Q}_{3}=\hat{A}^{{\dagger}}$), and so its eigenvalues $\lambda_{i}\in\mathbb{R}$ (i=0,1,,Ni=0,1,...,N) for a truncated $(N+1)$-dimensional subspace of the Hilbert space are real. The greatest eigenvalue $\lambda_{\text{max}}=\text{max}(\{\lambda_{i}\})$ corresponds to the largest possible Tsirelson violation, $\langle A\rangle>1$, for a superposition $|\psi\rangle=\sum_{n=0}^{N}c_{n}|n\rangle$ of QHO energy eigenstates up to $|N\rangle$. Likewise, the lowest eigenvalue corresponds to the most severe violation of the classical lower bound, $\langle A\rangle=-1$, in the subspace. The matrix elements of $A$ in the energy eigenbasis can be written in terms of the familiar $Q_{nk}$:
${A}_{nk}\equiv\langle n|{A}|k\rangle=Q_{nk}(1+e^{i\tau(n-k)}+e^{2i\tau(n-k)})\ ,$ (2.8)
for n,k=0,1,,Nn,k=0,1,...,N.
By explicitly diagonalising the Tsirelson operator $A$ in an increasingly large subspace, which is detailed in Appendix D, we find that the maximally-violating state in the $N=6$ subspace is given by:
$|\Psi_{\mathcal{Z}}\rangle=\frac{4}{\sqrt{42}}|0\rangle-\frac{1}{\sqrt{2}}|3\rangle+\sqrt{\frac{5}{42}}|6\rangle\ ,$ (2.9)
with a upper-bound Tsirelson violation $\langle A\rangle=1.1195$; or equivalently, by swapping the sign of the odd-parity eigenstate,
$|\Psi_{\mathcal{Z}}^{\prime}\rangle=\frac{4}{\sqrt{42}}|0\rangle+\frac{1}{\sqrt{2}}|3\rangle+\sqrt{\frac{5}{42}}|6\rangle\ ,$ (2.10)
with a lower-bound Tsirelson violation $\langle A\rangle=-1.1195$. This agrees exactly with the work of Ref. [6], for which we label Eq. 2.9 the Zaw–Scarani state. We shall use this as a simple candidate state for later analysis.
This is not an isolated violating state. Fig. 4 depicts a modified “Bloch sphere” representation of the set of all possible states constructed in the subspace spanned by the $|0\rangle$, $|3\rangle$, and $|6\rangle$ energy eigenstates with real coefficients. That is,
$|\Psi\rangle=a_{0}|0\rangle+c_{a}|3\rangle+a_{6}|6\rangle\ ,\ \text{where}\ a_{0},a_{3},a_{6}\in\mathbb{R}\ .$ (2.11)
Clearly, there exists a sizeable region of states over which a Tsirelson violation can be achieved to varying extent. This is good news for experimental realisations of the protocol, because it means even quantum states that are prepared imperfectly can, in principle, produce demonstrable Tsirelson violation. Another important observation is that every upper-bound-violating state is paired with a commensurate lower-bound-violating state ($A\rightarrow-A$), which is obtained by swapping the sign of the odd-parity coefficients (e.g. $c_{3}\rightarrow-c_{3}$).
Figure 4: Representation of all states in the subspace described by Eq. 2.11, where the axes represent the (real) coefficients. (a) A colour map is used to represent the value of the Tsirelson quantity when each of these states is taken to be the initial state. The Zaw–Scarani state, labelled $|\Psi_{Z}\rangle$, is the maximally violating state and lies in the centre of the upper-bound-violating region. (b) Only the violating (non-classical) regions are highlighted, showing that each upper violation is paired with a lower violation.
II.3 Structure of the Tsirelson operator
This section establishes a link between the Tsirelson operator and the group $C_{3}$. We then apply this connection to find an alternative explanation for a result previously noted in Ref. [4].
The group $C_{3}$ has three elements which we denote $\{e,r,r^{2}\}$. It is an Abelian group whose generator is $r$ [62]. From the Cayley table of the group, outlined in Table 1, it is straightforward to determine that $\{\mathds{1},U,U^{2}\}$ forms a faithful representation of $C_{3}$ under multiplication with $U$ acting as the generator.
$e$ $r$ $r^{2}$
$e$ $e$ $r$ $r^{2}$
$r$ $r$ $r^{2}$ $e$
$r^{2}$ $r^{2}$ $e$ $r$
Table 1: The Cayley table of $C_{3}$. The three elements of the group are $\{e,r,r^{2}\}$, where $e$ is the identity.
From this observation, we can relate the Tsirelson operator to the twirl of the sign operator,
$\operatorname{Twirl}(\hat{Q})=\frac{1}{3}A=\frac{1}{3}\left[\mathds{1}Q\mathds{1}+U^{\dagger}QU+(U^{2})^{\dagger}Q(U^{2})\right]\ .$ (2.12)
The $\operatorname{Twirl}(\cdot)$ operation is the averaging of a matrix over the actions of a group, in this case $C_{3}$ [63]. It is a useful operation because it extracts the component of the matrix that transforms under the identity (often dubbed the “trivial”) representation of the group. Note that $U^{\dagger}\operatorname{Twirl}(Q)U=(1)\operatorname{Twirl}(Q)\implies[\operatorname{Twirl}(Q),U]=0$.
Because the two normal operators commute, they cannot mix each other’s eigenspaces. Notice that the action of $U$ on the basis states $|n\rangle$ clearly partitions the Hilbert Space, $\mathcal{H}$, into 3 degenerate eigenspaces corresponding to its three eigenvalues $\lambda_{j}$:
Eigenspace for $\lambda_{1}=1$:
$\mathcal{H}_{0}=\operatorname{span}\{|n\rangle:n\equiv 0\mod 3\}\ .$ (2.13)
Eigenspace for $\lambda_{2}=e^{-i\frac{2\pi}{3}}$:
$\mathcal{H}_{1}=\operatorname{span}\{|n\rangle:n\equiv 1\mod 3\}\ .$ (2.14)
Eigenspace for $\lambda_{3}=e^{-i\frac{4\pi}{3}}$:
$\mathcal{H}_{2}=\operatorname{span}\{|n\rangle:n\equiv 2\mod 3\}\ .$ (2.15)
[…]
Figure 5: The Tsirelson operator in the energy eigenbasis with any finite energy cut-off can be interpreted as the adjacency matrix of a simple weighted graph. This is illustrated with a cut-off at $n=29$, where the nodes are energy eigenstates and the edges represent non-zero matrix elements of the Tsirelson operator (the colour representing the value of the matrix element).
III Assessing the Dynamical Assumption
III.1 General Framework
We now come to the main issue, which is assessing the degree to which the dynamical assumptions behind the Tsirelson inequality hold. As discussed, like many other classical phenomena, uniform precession and its closely-related assumptions have no unique definition in quantum theory. We will therefore explore a number of different, inequivalent implementations which entail different measurement techniques to check it. Specifically, we will implement the following procedure in a number of different ways:
  1. 1.
    Declare a data set, comprising measurements of $\langle A\rangle$ (the Tsirelson test itself) as well as subsidiary quantities to gauge the dynamical assumption of uniform precession (UP).
  2. 2.
    Write the Tsirelson quantity in the following form:
    $\frac{1}{2}(1+\langle A\rangle)=(\text{Positive term})+(\text{UP violating term})\\ +(\text{Quantum interference term})\ ,$ (3.1)
    where the scaled Tsirelson quantity $\frac{1}{2}(1+\langle A\rangle)$ is bounded between 0 and $1$ classically and hence describes a probability.
  3. 3.
    Establish whether or not the UP violating term is capable of dominating the quantum interference terms in terms of explaining the Tsirelson inequality violation. In other words, whether or not a classical model with no quantum-mechanical interferences could account for the violation. If not, the failure of macrorealism per se must also be implied.
III.2 Method 1: Macrorealistic Description of Sign Changes
We shall quantify the breaking of the dynamical assumption of uniform precession through the LG3 quantity $L_{1}$ (Eq. 1.12), corresponding to the quasi-probability sum q(+,+,+)+q(,,)q(+,+,+)+q(-,-,-), which should be zero classically.
Step 1: The full data set comprises all three correlators $C_{12},C_{23},C_{13})$, measured non-non-invasively, and $\langle A\rangle$ via a Tsirelson test.
Step 2: We find, using the probability moment expansions in Eqs. 1.7 and 1.10, that the Tsirelson quantity can be written as
$\frac{1}{2}(1+\langle A\rangle)=\sum_{j>i}p_{ij}(+,+)-L_{1}+I\ ,$ (3.2)
where the sum of two-time measurement probabilities,
$\sum_{j>i}p_{ij}\equiv p_{12}(+,+)+p_{23}(+,+)+p_{13}(+,+)\ ,$ (3.3)
is manifestly positive, and the interference term is given by
$I=\frac{1}{4}\left(\langle Q_{2}\rangle-\langle Q_{2}^{(1)}\rangle\right)+\frac{1}{4}\left(\langle Q_{3}\rangle-\langle Q_{3}^{(2)}\rangle\right)\\ +\frac{1}{4}\left(\langle Q_{3}\rangle-\langle Q_{3}^{(1)}\rangle\right)\ .$ (3.4)
This constitutes the explicit form of Eq. 3.1.
The candidate state we consider is the parity-flipped Zaw–Scarani state, Eq. 2.10. with a (re-scaled) lower-bound Tsirelson violation
$\frac{1}{2}(1+\langle A\rangle)=-0.0598<0\ ,$ (3.5)
which forms the left-hand side of the key formula (Eq. 3.2).
Step 3: Finally, we ask the question: is it possible that a non-zero no-crossing probability (that breaks uniform precession) explains the full Tsirelson violation? The first term on the right-hand side of Eq. 3.2 is positive, and we have calculated $\frac{1}{2}(1+\langle A\rangle)=-0.0598$ and $L_{1}=0.0421$ (from Appendix F). It is immediately clear that the violation of uniform precession is not sufficient to explain the Tsirelson inequality violation; hence non-zero quantum interferences are present. This is a crucial result, showing that this method has proven successful in identifying that uniform precession is satisfied well enough that quantum behaviour must be implied.
We also note that Eq. 3.2 can instead be written in terms of the two-time quasi-probabilities (Eq. 1.6), with no interferences:
$\frac{1}{2}(1+\langle A\rangle)=q_{12}(+,+)+q_{23}(+,+)+q_{13}(+,+)-L_{1}\ .$ (3.6)
Each of the two-time quasi-probabilities satisfies the the Lüders bound; but, in fact, the sum $q_{12}(+,+)+q_{23}(+,+)+q_{13}(+,+)$ is itself bounded below by $\frac{1}{8}$, as is readily shown by writing the sum purely in terms of $\langle A\rangle$ and $\langle A^{2}\rangle$ using the moment expansions (Eq. 1.6):
$\sum_{j>i}q_{ij}(+,+)=\frac{1}{8}\langle(A+2)^{2}-1\rangle\ .$ (3.7)
From Eq. 3.6 and Eq. 3.7, we see that the Tsirelson violation, $\frac{1}{2}(1+\langle A\rangle)$, will itself respect the Lüders bound if uniform precession is satisfied exactly ($L_{1}=0$), but in general can exceed this bound, as is the case with the maximally violating state in the full infinite-dimensional QHO Hilbert space, which Tsirelson found to be $\frac{1}{2}(1+\langle A\rangle)=-0.13<-0.125$ [1].
We now proceed with some further analysis on the breaking of uniform precession in this state. We already noted above that $L_{1}>0$, and we also found that $L_{2}=L_{3}=L_{4}=0.3193>0$. This complete satisfaction of the LG3s is not some esoteric feature unique to this simple candidate state we have chosen: in fact, we find in Appendix I that any Tsirelson-violating state constructed in one of the three subspaces given by Eqs. 2.132.15 will satisfy all four LG3 inequalities.
Recall from Eqs. 1.161.19 that the LG3 quantities describe a probability distribution of sign changes. Therefore, there exists a valid classical (macrorealistic) description of crossing probabilities for the state. That there exists a macrorealistic description of sign-change probabilities is a crucial point: we now have the basis to claim that $L_{1}$ represents a valid probabilistic description of the breaking of uniform precession.
The interference term in Eq. 3.1 is purely quantum-mechanical in nature; however, an important problem to consider is whether any further non-classicality is indicated by $L_{1}$, which should be zero classically. Non-classicality in temporal correlations can stem from two separate sources: non-commutativity of successive measurements and negativity of the Wigner function [64]. Whilst violations of the Tsirelson inequality require negativity of the Wigner function, the non-zero positive $L_{1}$ is an independent diffraction-in-time effect, as characterised in Ref. [65] but first studied in Refs. [66, 67]. Therefore, even a state with positive Wigner function, which will not violate the Tsirelson quantity, can render this LG3 non-zero.
More speculatively, we also note that the size of the uniform precession violation here is roughly $4\%$, which is peculiarly similar to the estimated bound on the so-called quantum back-flow effect, the phenomenon in which a particle prepared with strictly positive momentum can still exhibit negative probability current [68, 69, 17, 70].
We also note, at this point, that simple superpositions of coherent states—such as the “three-headed cat state”—will violate the Tsirelson quantity [5]. We have also verified in our calculations that the state
$|\Psi\rangle=N\left(|\sqrt{2}\ \rangle+|e^{2\pi i/3}\sqrt{2}\ \rangle+|e^{4\pi i/3}\sqrt{2}\ \rangle\right)\ ,$ (3.8)
with the appropriate normalisation factor $N$, yields a Tsirelson violation of $\langle A\rangle=-1.0795<-1$. Thus, the method of assessing UP with an LG quantity can also be used upon such states, which may be of interest experimentally, for Gaussian state superpositions can be easier to create in a laboratory setting [71].
III.3 Method 2: Three-time Sequential Measurements
Our second method uses a generalised dynamical assumption involving $\Delta p=p(+,+,+)-p(-,-,-)$, where $p$ is a probability to be specified. This is of interest because, classically, the Tsirelson quantity $A$ can take values $\pm 3$ and $\pm 1$, but setting $\Delta p=0$ prevents the values $\pm 3$ from contributing to the computation of the average $\langle A\rangle$, and hence trivially $1\leq\langle A\rangle\leq 1$.
The probabilities could be specified to involve three-time projective measurements, $\Delta p_{123}$, corresponding to Eq. 1.11, or “weak” measurements implemented at times $t_{1}$ or $t_{2}$, denoted by $\Delta p_{(1)23}$ and $\Delta p_{1(2)3}$, respectively. Finally, we could also consider the quasi-probability $\Delta q$, corresponding to Eq. 1.7. Appendix H details the computational approach to calculating all such measurement probabilities. An important point is that all of these four distributions are identical to a macrorealist but differ quantum mechanically by various interference terms.
Step 1: We take the data set to be $\Delta p$ measured via various three-time measurements and the Tsirelson quantity $\langle A\rangle$ measured with single-time measurements.
Step 2: We find an alternative way in which the Tsirelson quantity can be expressed, using this time the moment expansion Eq. 1.11:
$\frac{1}{2}(1+\langle A\rangle)=\frac{1}{2}(1-D)+2\Delta p+I\ .$ (3.9)
Here $I$ is the interference term, and its values for each of the four cases are given below in Table 2. Both classically and quantum-mechanically, the triple correlator satisfies $|D|\leq 1$, which makes the first term $\frac{1}{2}(1-D)$ manifestly non-negative, so that Eq. 3.9 is the explicit form of Eq. 3.1 in this method.
Step 3: We compare the size of the UP violation, $2|\Delta p|$, with the size of the Tsirelson violation $\frac{1}{2}|(1+\langle A\rangle)|$, noting once again that the interference terms would vanish in a purely classical model of the violation.
Table 2 catalogues our numerical results with this approach, considering each of the four aforementioned distributions. We compare the extent of a lower-bound violation that would be possible from only the breaking of the dynamical assumption in a classical model with the true value of the Tsirelson violation, using the same state as before (Eq. 2.10). With the three-time sequential measurement probability $\Delta p_{123}$ as the chosen measure of UP, we calculate that $2|\Delta p_{123}|=0.0191<0.0598$, showing once again that the breaking of the dynamical assumption falls significantly short of accounting for the full extent of the Tsirelson violation, thereby implying the presence of quantum-mechanical interference term.
When weak measurements are employed, UP is satisfied only marginally in the case of $\Delta p_{1(2)3}$; and when the quasi-probability distribution $\Delta q$ is instead used, which produces no interference terms, the UP violation is exactly commensurate with the Tsirelson violation. This is to be expected, since $\Delta q=0$ is a sufficient condition to derive the Tsirelson inequality, which follows directly from the moment expansion form of Eq. 1.7. The quasi-probabilities, though they can be negative, respect the sum rules for probabilities; when a measurement is rendered weak, the probability sum rule holds valid over that particular measurement time, as in $p_{1(2)3}(+,+,+)+p_{1(2)3}(+,-,+)=p_{13}(+,+)$.
  Measure of UP
Interference term, II
$2\Delta p$
UP satisfied?
($2|\Delta p|<\frac{1}{2}|1+\langle A\rangle|=0.0598$ ?)
$\Delta q$ 0 0.0598-0.0598
$\Delta p_{123}$
$\frac{1}{2}(\langle\hat{Q}_{2}\rangle-\langle\hat{Q}_{2}^{(1)}\rangle)+\frac{1}{2}(\langle\hat{Q}_{3}\rangle-\langle\hat{Q}_{3}^{(12)}\rangle)$
0.0191-0.0191
$\Delta p_{(1)23}$ $\frac{1}{2}(\langle\hat{Q}_{3}\rangle-\langle\hat{Q}_{3}^{(2)}\rangle)$ 0.0785-0.0785
$\Delta p_{1(2)3}$
$\frac{1}{2}(\langle\hat{Q}_{2}\rangle-\langle\hat{Q}_{2}^{(1)}\rangle)+\frac{1}{2}(\langle\hat{Q}_{3}\rangle-\langle\hat{Q}_{3}^{(1)}\rangle)$
0.0573-0.0573
Table 2: The various means by which the dynamical assumption of uniform precession (UP) can be assessed through three-time probability distributions in Method 2 of Section III. A subscript in brackets indicates that the measurement at that time was weak. A tick indicates that the UP violation is too small to produce the overall violation, which must therefore be produced by the quantum-mechanical interference terms.
An alluring question is whether it is possible, by some variation of the parameters for the quantum state, to render $\Delta p_{123}=0$ exactly, whilst remaining within a Tsirelson violating region (such as that depicted in Fig. 4), which may serve the poetic justice of “total satisfaction” of the dynamical assumption, whereby the Tsirelson violation is solely due to effects quantum-mechanical in origin. In fact—unlike in Method 1, wherein setting $L_{1}=0$ leads trivially to the satisfaction of the Tsirelson inequality—it is possible in this case to set $\Delta p_{123}=0$. We find by using a constrained optimisation procedure via Lagrange multipliers that a simple state, still within the $N=6$ subspace, albeit now with complex coefficients, can be found to produce $\langle A\rangle=1.0756$ with $\Delta p_{123}=0$. This procedure is documented in Appendix J.
IV Heuristic Notions of Uniform Precession
In this section, we describe an alternative method of assessing uniform precession which is somewhat heuristic and indirect but wherein the measurements involved are simpler and in some cases are, appealingly, single-time measurements. We consider measurements of two simple quantities during a given time interval: the dwell time in $x>0$ and the number of crossings of the origin. Physical measurements will reveal the quantum-mechanical values of these quantities and, we shall show, if sufficiently close to the corresponding classical values will significantly constrain possible violations of uniform precession in a trajectory model. This in turn gives heuristic indications of the smallness of the various forms of the probabilities $p(+,+,+)$ and p(,,)p(-,-,-) used in Methods 1 and 2 above. We also relate the Tsirelson inequality to the probability current.
IV.1 Dwell time
We mentioned previously that dynamical assumptions are often classically motivated. One potential notion of how well the dynamical assumption of uniform precession is satisfied is the “dwell time”, which is classically given by the following integral:
$T_{D}(\tau)=\frac{1}{2\pi}\int_{0}^{\tau}\theta(x(t))dt\ .$ (4.1)
The dwell time is a measure of the fraction of time the particle spends in $x>0$. Classically, this value should be $1/2$ over a full period ($\tau=2\pi$).
This can readily be turned into a self-adjoint operator, $\hat{T}_{D}(\tau)$ [72], and we can take its expectation value, $\langle\hat{T}_{D}(\tau)\rangle$. This quantum-mechanical dwell time can be ascertained experimentally by measuring $\langle\hat{\theta}(x(t))\rangle$ at multiple times.
For an arbitrary state $|\psi\rangle=\sum_{n}c_{n}|n\rangle$, in terms of energy eigenstates, we can write
$\langle\hat{T}_{D}(2\pi)\rangle=\sum_{n}|c_{n}|^{2}\langle n|\hat{\theta}(t)|n\rangle\ .$ (4.2)
As $|n\rangle$ are stationary states, $\langle n|\hat{\theta}(t)|n\rangle=\langle n|\hat{\theta}(0)|n\rangle$, and because energy eigenstates have definite parity, $\langle n|\hat{\theta}(0)|n\rangle=\frac{1}{2}$. Therefore,
$\langle\hat{T}_{D}(2\pi)\rangle=\frac{1}{2}.$ (4.3)
We see that the dwell time, integrated over a full period, is exactly the classical value. [^2] In fact, the dwell time operator in the energy eigenbasis is proportional to the identity. By this measure, a classical and quantum oscillator are indistinguishable.
[…]
It is quite striking that the quantum dwell time agrees with the classical result and yet Tsirelson violations still occur. In fact, Ref. [4] considered alternatives to the Tsirelson scheme with a greater number of measurement times during a period and found that quantum violations faded away as the number of measurement times was increased (limiting to the dwell time result which has no quantum violations). This perspective motivates viewing the Tsirelson quantity as a Riemann sum approximation to the dwell time integral. Upon performing such an analysis (see Appendix K), we have found that the Tsirelson quantity is related to the probability density current, $J(x,t)=\operatorname{Im}\left(\psi^{*}\partial_{x}\psi\right)$, at the origin:
$\langle A\rangle=(6\langle\hat{T}_{D}(2\pi)\rangle-3)+2\pi J(0,\tau)=2\pi J(0,\tau)\ ,$ (4.4)
where $\tau$ is some time on the interval $0,2\pi]$, as the result was obtained using the mean value theorem. Thus, we see that the dependence of $\langle A\rangle$ upon the average dwell time drops out, since $\langle\hat{T}(2\pi)\rangle$ assumes the classical value, and the Tsirelson violation is fundamentally rooted in the probability current at the origin.
IV.2 Crossing number
Another simple property of classical uniform precession, which is closely linked to Tsirelson’s dynamical assumption, is that trajectories cross the origin exactly twice over a period (barring the singular case of a particle sitting at rest at the phase-space origin). Formally, we can compute crossings over a time interval $\tau$ using
$N_{c}(\tau)=\int_{0}^{\tau}\mathrm{d}t|p(t)|\delta(x(t))\ ,$ (4.5)
where $N_{c}(2\pi)=2$.
Motivated by this classical expression, we define the Hermitian quantum operator
[…]
where $\operatorname{sign}(\hat{p}(t))$ and $\delta(\hat{x}(t))$ are defined via the functions being applied to the eigenvalues of the spectral decomposition. We can expand out $\hat{N}_{c}(\tau)$ as
$\hat{N}_{c}(\tau)=\frac{1}{4}\int_{0}^{\tau}\mathrm{d}t\iint\mathrm{d}p\ \mathrm{d}q\operatorname{sign}(p)qe^{iHt}\left[A+B\right]e^{-iHt}\ ,$ (4.7)
where
$A=|p\rangle\langle p|q\rangle\langle q|x_{0}\rangle\langle x_{0}|+|x_{0}\rangle\langle x_{0}|q\rangle\langle q|p\rangle\langle p|\ ,$ (4.8)
$B=|q\rangle\langle q|x_{0}\rangle\langle x_{0}|p\rangle\langle p|+|p\rangle\langle p|x_{0}\rangle\langle x_{0}|q\rangle\langle q|,$ (4.9)
and where $\langle x|p\rangle=\frac{1}{\sqrt{2\pi}}e^{ipx}$ and $|x_{0}\rangle$ is an eigenstate of the position operator at the origin. Then, when $\tau=2\pi$, the operator is diagonal in an energy eigenbasis and the diagonal elements are
$\langle n|\hat{N}_{c}(2\pi)|n\rangle=\frac{\pi}{2}\iint\mathrm{d}p\ \mathrm{d}q\operatorname{sign}(p)q\left[\langle n|A|n\rangle+\langle n|B|n\rangle\right]\ .$ (4.10)
But more conveniently, when Eq. 4.6 is written as
[…]
we note that the first term picks up only symmetric states and the second term only anti-symmetric states.
We observe an asymptotic tendency towards the classical value of two crossings in the limit of high energy eigenstates, which is depicted in Fig. 6. Moreover, we calculate for the Zaw–Scarani state an average crossing number of $\langle\hat{N}_{c}(2\pi)\rangle_{Z}\approx 1.49$, with a standard deviation of $\Delta\hat{N}_{c}(2\pi)_{Z}\approx 0.2$. Over a third of a period, between measurement times, $\hat{N}_{c}$ remains diagonal for states in $\mathcal{H}_{0}$, and so $\langle\hat{N}_{c}(2\pi/3)\rangle_{Z}\approx 0.50$ with a standard deviation of $\Delta\hat{N}_{c}(2\pi)_{Z}\approx 0.07$. This means that the probabilities $p(+,+,+)$ and p(,,)p(-,-,-), which underpin the assumption of uniform precession in the Tsirelson scheme, must be very small indeed, since this would require either no crossings at all or at least two crossings between any pair of measurement times. The spread in crossings over a third of a period suggests that this situation is statistically very unlikely; nonetheless, we have not found an exact connection to the relevant probabilities.
Figure 6: Plot of the average crossing number, $\langle n|\hat{N}_{c}(2\pi)|n\rangle$ (see Eq. 4.10), over a full period for the first 21 energy eigenstates. We see that the even eigenstates converge much more rapidly to the classical value of $2$.
A potential issue is that measuring the crossing number requires a two-time measurement, which could raise concerns of invasiveness once again. However, we note that, classically,
$|p|\delta(x)=\sqrt{p^{2}+x^{2}}\delta(x)=\sqrt{2E}\delta(x)$ (4.12)
in a fixed energy state. This in turn means that a semi-classical approximation can be made for symmetric states, wherein the first term of Eq. 4.11 can be taken to be $\sim\sqrt{2E}|\psi(0)|^{2}$. Via this approximation, only single-time measurements are required. However, states like the Zaw–Scarani contain a large anti-symmetric part; also, the spread $\Delta\hat{N}_{c}$ cannot be determined with single-time measurements, and thus this approach is only partially successful.
IV.3 Implications for trajectory models
We have seen that the quantum-mechanical properties of the dwell time are identical to its classical properties and that as for the crossing number, its quantum-mechanical properties are qualitatively the same as classical and quantitatively similar. These features suggest that measurements of the dwell time and crossing number may be used to give indirect but useful bounds on the size of uniform precession violations. These quantities are, in principle, related to the three-time probabilities used in Section III, but any such relation will be non-unique, and we have not been able to find a particularly useful one. Therefore, we proceed instead using a simpler, more heuristic analysis.
We will suppose that there exists an underlying probability distribution on a set of phase-space trajectories $p(t)$, $x(t)$ which are close to classical trajectories but could potentially violate the Tsirelson inequality by violating uniform precession. We then examine what restrictions on these trajectories are implied by measurements of the dwell time and crossing number, with a view to ruling out these alternative classical explanations of the Tsirelson violations.
The properties of the dwell time, Eq. 4.3, over the time interval $0,2\pi]$ indicate that the underlying trajectories must each spend equal amounts of time in $x>0$ and $x<0$. However, as noted, the dwell time must have the same properties over any shifted time interval of the same size. The implication in phase space is the trajectories must spend equal times above or below any infinite radial line. Thus, the set of trajectories must have rotational symmetry, which suggests that the only possibility is a set of circles around the origin with constant angular velocity along each circle. This still allows a set of circular trajectories with a spread of possible constant angular velocities, which will cause a violation of uniform precession. However, trajectories with angular velocities faster or slower than the classical one for the given Hamiltonian will create non-classical values for the crossing number. But as the crossing number results above show, and as noted, the probabilities for four or more crossings or no crossings, which would be responsible for violations of uniform precession, are in fact very small.
IV.4 Classical versus quantum dynamics
Finally, we make some general remarks about dynamics in classical versus quantum theories and why we might expect, on general grounds, that the violation of uniform precession be small. The predictions of quantum theory are very different to the predictions of classical physics, and most of these differences trace back to the simple fact that quantum theory allows a much richer variety of states than does classical physics—for example, through Wigner functions that can be negative versus classical phase-space probabilities. This is responsible for some of the most striking quantum phenomena, such as Bell inequality violations.
However, classical and quantum physics look at lot more similar at the level of dynamics, and there are many results in the history of quantum theory that prove this. For example, wave packets—or more generally, phase-space localised quasi-projectors—evolve to a good approximation along classical paths (see for example Ref. [73]). This in turn is related to the fact that the Wigner function, for a class of potentials, has approximately classical evolution. This behaviour can be affected by monitoring (e.g. with projection operators onto ranges of position at intermittent times), but this effect is not large unless the projections are onto very small regions and the projectors are applied very frequently, which is determined by the energy time $\hbar/\langle E\rangle$ [74]. (For the Zaw–Scarani state considered in this paper, the energy time is calculated to be $1/12$ of a period.) Furthermore, specific models, such as LG2 violations with a Gaussian initial state, produce some non-classicality, but only about $30\%$ of the level of non-classicality that can be produced by a negative Wigner function [64, 65]. Of course, this discussion involves quantum-mechanical properties. However, the sceptical macrorealist could confirm these ideas using control experiments in which the dynamics of a system in a set of initially classical-like states is measured.
So, in summary, there are general physical and experimentally-testable reasons why we might expect the violation of a dynamical assumption like uniform precession to be reasonably small compared to the more significant non-classical effects that can arise from the initial state.
V Summary and Discussion
It is remarkable that a simple assumption about the classical dynamics of an oscillator—so straightforward that one would think it must surely be upheld universally—in fact reveals ever more of the bizarre and distinguishing predictions of quantum theory. The Tsirelson inequality also makes clear that, by introducing a dynamical assumption, the prospect of observing non-classicality in temporal correlations need not be plagued by the problem of circumventing invasive sequential measurements.
In Section II, we analysed maximally-violating states and formalised the Tsirelson operator as the twirl of the sign operator over the group elements of the three-element cyclic group. This perspective directly explained its eigenspace structure and opens up the possibility for more tools from quantum information to provide further insight into the Tsirelson operator. We provided proofs that a superposition of at least three energy eigenstates of the QHO is required to produce a Tsirelson violation and that, in relevant subspaces, all three correlators are equal for the specific measurement time interval of two-thirds of a period. The violations possible with simple states such as the Zaw-Scarani state—and, moreover, superpositions of coherent states—are certainly promising and within the reach of experimentalists in the coming years.
But, as we noted in Section I, there is a crucial loophole in the Tsirelson inequality: particularly, that violations could ensue from the failure of the dynamical assumption of uniform precession in a classical system, meaning that a Tsirelson violation is not a sufficient condition for genuine quantum behaviour. Therefore, in this paper we have developed a subsidiary protocol to disentangle violations of macrorealism from the breaking of the dynamical assumption—to which end we have embedded the Tsirelson inequality within the wider Leggett–Garg formalism.
In Section III, we broke apart the Tsirelson quantity into separate terms that quantified the breaking of uniform precession and quantum interferences, such that their sizes may be compared. Recognising that uniform precession can be defined quantum-mechanically in several inequivalent ways in terms of measurable probabilities, we investigated two methods. In the first method, we proved the surprising fact that eigenstates of the Tsirelson operator satisfy all four LG3 inequalities, meaning that there invariably exists a macrorealistic description of sign changes, which can then be used to assess uniform precession through an LG3 test. In the second method, we considered a generalised dynamical assumption with the difference of no-sign-change probabilities (perhaps less desirable than the first method, for it involves three sequential measurements). In all cases, we found success with the protocol, demonstrating that quantum interference terms must be implied, elevating the Tsirelson test from a simple test of non-classical behaviour to a more robust demonstration of quantumness.
In Section IV, we presented further qualitative evidence that uniform precession is satisfied well in states that violate the Tsirelson inequality, by considering the heuristic notions of dwell times and the crossing number operator. We found that the quantum-mechanical dwell time exactly obeys the classical bound, and that the total number of origin crossings is approximately the classical value. On their own, these stand as interesting results, further elucidating the various ways in which uniform precession is and is not equivalent in the classical and quantum cases. But we also further interpreted these quantities in the context of underlying trajectory models, noting that they greatly restrict the extent to which underlying trajectories can be erratic. We also came upon the global property that Tsirelson violations are linked to the probability currents through the origin, whereby weakly-flowing states, or the classicalised currents of quantum states, never produce a violation.
We conclude with a discussion on the key issue of whether or not the need for sequential measurements has really been eliminated—a virtue of the Tsirelson inequality we so admired in the first place—if, at the very least, a Leggett–Garg test is still required to assess uniform precession. Is non-invasive measurement then re-invoked as an untoward assumption alongside macrorealism per se and uniform precession? At one level, this paper can be regarded as a purely quantum-mechanical analysis of Tsirelson violations, and we have found that they arise largely from the presence of interferences and a genuine failure of macrorealism, but with a small contribution stemming from UP violations.
However, an ultimate goal for Tsirelson tests will be to probe systems not knowing beforehand whether they are quantum or not. Then, to an experimentalist—who should wish to implement this protocol by collecting a data set of the Tsirelson quantity supplemented with a Leggett–Garg test for the relevant probabilities—the problem of re-invoking a Leggett–Garg test, in order that we should further conclude that macrorealism per se has been violated, is possibly harder to justify. But we may appeal here to the heuristic results of Section IV, which provide evidence that the LG3 quantity $L_{1}$ should be small and which support the idea that violations of uniform precession are not, in general, sufficiently large to be the primary cause of Tsirelson inequality violations. In some cases, we were partially successful in showing how these quantities may be measured through only single-time measurements. In the original Leggett–Garg scheme, there is no reason to expect a priori that $L_{1}$ should be small; here, however, we have gathered evidence that it should indeed be small (and positive), and so if an experiment finds otherwise, there is reason to suspect that measurement invasiveness may be largely to blame.
The results of this paper open up several new research directions. Future work could consider an anharmonic oscillator, multiple coupled oscillators and links to field theories, entanglement, alternative measures of crossings, and stronger quantum-mechanical definitions of uniform precession, amongst other things. We have concentrated in this work upon the case of the harmonic oscillator, but it would be interesting to conduct this same analysis and assessment of uniform precession for spin systems; subsequent comparison with the results herein obtained could be insightful. The idealistic and experimentally impracticable nature of perfect sign operator projective measurements could be remedied by extending this analysis to use smoothed projectors. The analysis of Tsirelson’s inequality and the breaking of uniform precession could be extended to incorporate suspected connections with the quantum back-flow limit. This could all provide further illumination on the true nature of violations of Tsirelson’s intriguing inequality and, altogether, bring us closer to loophole-free tests of quantum behaviour.
Acknowledgements.
We are grateful to Clement Mawby for useful discussions and Lin Htoo Zaw for bringing to our attention a result regarding superpositions of coherent states.
Appendix A Tsirelson Quantity from the Wigner Function
The Wigner quasiprobability distribution $W_{\rho}(x,p)$ is given by
$W_{\rho}(x,p)=\frac{1}{2\pi i}\text{Tr}\left(\rho\ e^{\frac{i\pi}{2}\left[(\hat{X}-x)^{2}+(\hat{P}-p)^{2}\right]}\right)\ .$ (A1)
Thus, the Tsirelson quantity can also be written as
$\langle A\rangle=\text{Tr}(\rho A)=\int\int dx\ dp\ W_{A}(x,p)W_{\rho}(x,p)\ ,$ (A2)
where $W_{A}(x,p)$ is the phase-space representation of the operator $A$, obtained from its Weyl transform:
$W_{A}(x,p)=\int_{-\infty}^{\infty}dy\,e^{-ipy/\hbar}\left\langle x+\frac{y}{2}\middle|A\middle|x-\frac{y}{2}\right\rangle\ .$ (A3)
This formula is valid for any Hermitian operator.
Appendix B Single crossing dynamical assumption
An alternative set of inequalities can be formulated under the single crossing assumption (SCA), which is an alternative dynamical assumption that instead asserts that p(+,,+)=p(,+,)=0p(+,-,+)=p(-,+,-)=0, forbidding that the particle cross the origin more than once. A very rapidly oscillating system could cross many times in between measurements whilst maintaining p(+,,+)=p(,+,)=0p(+,-,+)=p(-,+,-)=0, but we assume that the relative amplitudes of these erratic trajectories are small.
SCA can be imposed in the same way as was done uniform precession in the original Tsirelson derivation, by setting p(+,,+)=p(,+,)=0p(+,-,+)=p(-,+,-)=0. Using the moment expansions 1.7, it is straightforward to obtain
$1\leq\langle Q_{1}\rangle-\langle Q_{2}\rangle+\langle Q_{3}\rangle\equiv\langle A_{\text{SCA}}\rangle\leq 1\ .$ (B1)
One of the most interesting features of the original Tsirelson inequality, we found, was that it could be derived from the weaker condition that p(+,+,+)=p(,,)p(+,+,+)=p(-,-,-). In fact, there is an exact analogue of this for the SCA, whereby Eq. B1 follows directly from the weaker condition p(+,,+)=p(,+,)p(+,-,+)=p(-,+,-).
Classically, one of the simplest examples of a system for which the SCA is valid would be a free particle. Acted upon by no external forces, a free particle will only ever cross a pre-defined origin at most once. In the case of a uniformly-precessing system, such as the harmonic oscillator, this assumption holds true over half a period. We choose measurement times $t_{i}=\frac{iT}{6}$, $i=0,1,2$, where $T$ is the time period of the oscillator.
We repeat the same analysis as before, defining an operator
$A_{\text{SCA}}=\hat{Q}_{t=0}-\hat{Q}_{t=1/6}+\hat{Q}_{t=1/3}\ ,$ (B2)
where we have set the time period to unity for convenience. What we will now proceed to show is that imposing SCA over half a period is mathematically identical to imposing uniform precession over a full period, in that
$\langle A_{\text{SCA}}\rangle=\langle A\rangle\ .$ (B3)
Since $\langle A\rangle=\langle\hat{Q}_{t=0}\rangle+\langle\hat{Q}_{t=1/3}\rangle+\langle\hat{Q}_{t=2/3}\rangle$ and $\langle A_{\text{SCA}}\rangle=\langle\hat{Q}_{t=0}\rangle-\langle\hat{Q}_{t=1/6}\rangle+\langle\hat{Q}_{t=1/3}\rangle$, we see that these two quantities only differ by one term. If indeed it is true that $\langle A\rangle=\langle A_{\text{SCA}}\rangle$ then it is implied that $\langle\hat{Q}_{t=2/3}\rangle=-\langle\hat{Q}_{t=1/6}\rangle$. In fact, the more general result holds true that $\langle\hat{Q}_{t}\rangle=-\langle\hat{Q}_{t+1/2}\rangle\ \forall\ t$. That is to say, the single-time average of the dichotomous variable $Q$ merely flips sign after half a period.
We can understand why this is the case by considering what happens to the energy eigenfunctions after exactly half a period. Undergoing unitary time evolution in accordance with the QHO Hamiltonian, the eigenstate $|n\rangle$ accrues a relative phase of $e^{-i\pi n}$ after half a period. As such, the odd-$n$ eigenstates will pick up a minus sign, whilst the even-$n$ states are unaffected. In the language of wavefunctions, it is simple to see that this means a reflection across the $x$-axis after exactly half a period. Therefore, quantum mechanically, the quantities $\langle A\rangle$ and $\langle A_{\text{SCA}}\rangle$ are identical.
Appendix C Matrix Elements in an Energy Eigenbasis
Let
$|\Psi\rangle=\sum_{n=0}^{\infty}c_{n}|n\rangle,\qquad\hat{Q}=2\hat{\theta}-\mathds{1},$ (C1)
where $\hat{\theta}$ is the Heaviside projector on $x>0$. To obtain explicit matrices we truncate the Hilbert space at a finite excitation $N$:
$|\Psi\rangle=\sum_{n=0}^{N}c_{n}|n\rangle,\qquad\mathbf{c}=(c_{0},\ldots,c_{N})^{\!\mathsf{T}}.$ (C2)
For any operator $\hat{O}$ its expectation value reduces to a quadratic form,
$\langle\Psi|\hat{O}|\Psi\rangle=\sum_{n,k=0}^{N}c_{n}^{*}\,O_{nk}\,c_{k}=\mathbf{c}^{\dagger}\,O\,\mathbf{c},\qquad O_{nk}:=\langle n|\hat{O}|k\rangle.$ (C3)
Hence the task is to evaluate the matrix $Q_{nk}=\langle n|\hat{Q}|k\rangle$. Using the position representation and energy eigenstate wavefunctions $\Psi_{n}(x)$,
$Q_{nk}=\!\int_{-\infty}^{\infty}\!\!\Psi_{n}(x)\Psi_{k}(x)\,\bigl{[}2\theta(x)-1\bigr{]}\,\mathrm{d}x\ ,$ (C4)
using the property that the energy eigenstates for the 1D harmonic oscillator can always be taken to be real to drop the complex conjugate on $\Psi_{n}(x)$. To make the integral tractable, we split the integral at the origin and use the overlap formula in the appendix of Ref.[64],
$F_{nk}\equiv\int_{0}^{\infty}\Psi_{n}(x)\Psi_{k}(x)\,\mathrm{d}x=\frac{\Psi_{k}^{\prime}(0)\Psi_{n}(0)-\Psi_{n}^{\prime}(0)\Psi_{k}(0)}{2(k-n)}\quad(n\neq k),$ (C5)
where $F_{nk}=\langle n|\hat{\theta}|k\rangle$ and $F_{nn}=\frac{1}{2}$. Therefore, we find
$Q_{nk}=2F_{nk},\qquad Q_{nn}=0.$ (C6)
To evaluate Equation C5, we use the closed form expressions for the eigenstates of the harmonic oscillator in terms of the Hermite polynomials $H_{n}(x)$,
$\Psi_{n}(x)=\frac{\pi^{-\frac{1}{4}}}{\sqrt{2^{n}n!}}\exp\left(-\frac{1}{2}x^{2}\right)H_{n}(x).$ (C7)
Consequently, using the following identities for the derivatives of Hermite polynomials,
$H^{\prime}_{0}(0)=0,\qquad H_{n}^{\prime}(0)=2n\,H_{n-1}(0)\quad(n\geq 1),$ (C8)
we find
$Q_{nk}=\frac{2\pi^{-1/2}}{(k-n)\sqrt{2^{\,n+k}n!\,k!}}\bigl{[}k\,H_{k-1}(0)H_{n}(0)-n\,H_{n-1}(0)H_{k}(0)\bigr{]}\quad(n\neq k).$ (C9)
Finally, since $\hat{U}|n\rangle=e^{-2\pi in/3}|n\rangle$,
$A_{nk}=Q_{nk}\Bigl{[}1+e^{\tfrac{2\pi i}{3}(n-k)}+e^{\tfrac{4\pi i}{3}(n-k)}\Bigr{]}.$ (C10)
Our numerical implementation of the sign operator and Tsirelson operator on sympy, the symbolic python maths package, also employed the following identities:
$H_{2m}(0)=(-2)^{m}(2m-1)!!,\quad H_{2m+1}(0)=0.$ (C11)
We verified this implementation by checking matrix elements of the $F_{nk}$ agreed with those cited in Ref.[64] and our implementation of the Tsirelson matrix was verified by checking some known violating states.
Appendix D Diagonalisation of the Tsirelson operator in an energy eigenbasis
Note that when $\tau=\frac{2\pi}{3}$, the term
$1+e^{i\tau(n-k)}+e^{2i\tau(n-k)}=1+\cos(\frac{2\pi}{3}(n-k))+\cos(\frac{4\pi}{3}(n-k))+i\sin(\frac{2\pi}{3}(n-k))+i\sin(\frac{4\pi}{3}(n-k))$ (D1)
in Eq. 2.8 vanishes unless the difference (nk)(n-k) is a multiple of 3. Moreover, not only is the Zaw–Scarani state the maximally violating state in the $N=6$ subspace, but also in the $N=7$ and $N=8$ subspaces. For $N=9$, on the other hand, we can achieve a superior violation of $\langle A\rangle\approx 1.12$ with the state
$|\Psi_{9}\rangle\approx 0.6277|0\rangle-0.7064|3\rangle+0.3255|6\rangle+0.0304|9\rangle\ .$ (D2)
We resort to an approximate numerical specification of the coefficients, because neat analytic forms are intractable. Importantly, any numerical errors are encapsulated in the floating point specification of these coefficients, and the remaining coefficients ($c_{1},c_{2},c_{4},\text{etc...}$) really are exactly and identically zero.
Fig. 7 shows the results of a spectral decomposition of the Tsirelson operator in a high-energy truncated subspace. This serves as a good illustration that the classical values $A=\pm 1$ are “accumulation points”, but violations past this value are indeed possible. None of our computed eigenvalues exceed $|A|\approx 1.26$, which agrees with the value of the maximum possible violation quoted in Tsirelson’s original paper [1]. A more accurate estimate of the true upper and lower bounds on the spectral norm of $A$ has been obtained in the recent work of Ref. [9]. Fig. 8 illustrates that considering a progressively larger subspace from the low-lying energy states upwards does indeed lead to an asymptotic convergence towards Tsirelson’s quoted value (up to an arbitrary scaling factor) for the maximum violation in the full QHO Hilbert space.
Figure 7: A histogram of the spectrum of eigenvalues of $A$ is plotted by using an energy truncation of $E=70(\hbar\omega)$. The y-axis refers to the number of eigenvalues found within a given range.
Figure 8: The computed maximal eigenvalue of A in an increasingly large subspace, tending towards a finite maximal violation. Note, importantly, that we re-scale the eigenvalues as $\frac{1}{6}(A+3)$. This is only an arbitrary scaling choice to see agreement with the definition of the operator considered by Tsirelson in Ref. [2]. The Zaw–Scarani state ($|\Psi_{Z}\rangle$) is pointed out at $N=6$.
We have also computed the eigenvectors of $A$ in increasingly larger subspaces, catalogued in Table 3, and have found that they all obey the same pattern of being composed solely of energy eigenstates that are multiples of three.
  NN $c_{0}$ $c_{3}$ $c_{6}$ $c_{9}$ $c_{12}$ $c_{15}$ $c_{18}$ $\langle A\rangle$
3 $\frac{1}{\sqrt{2}}$ $\frac{1}{\sqrt{2}}$ 0 0 0 0 0 0.9772
6 $\frac{4}{\sqrt{42}}$ $\frac{1}{\sqrt{2}}$ $\sqrt{\frac{5}{42}}$ 0 0 0 0 1.1195
9 0.6277 -0.7064 0.3255 0.0304 0 0 0 1.1200
12 -0.6529 0.6600 -0.1591 -0.2538 0.2200 0 0 1.1500
15 -0.6509 0.6591 -0.1592 -0.2561 0.2259 -0.0083 0 1.1500
18 -0.6655 0.6354 -0.0999 -0.2771 0.1451 0.1396 -0.1615 1.1661
90 $|\Psi\rangle\rightarrow$ True eigenstate of $\hat{A}$ 1.2131
Table 3: Explicit computation of maximally violating states in a truncated energy subspace ($|\Psi_{N}\rangle=\sum_{n=0}^{N}c_{n}|n\rangle$). The only non-zero coefficients are $c_{3j}$, where $j\in\mathbb{N}$.
Appendix E Superpositions of Two Energy Eigenstates
As established in Ref.[1], single energy eigenstates of the harmonic oscillator cannot violate Tsirelson’s inequality due to their definite parity. The next logical step is to find violations produced by a superposition of two eigenstates. However, as we show now, any superposition of two energy eigenstates is insufficient to produce a violation.
E.1 Proof Strategy
We aim to show that the magnitude of $\langle A\rangle$ is less than $1$ for the state
$|\psi_{2}\rangle=\alpha|n\rangle+\beta|k\rangle,$ (E1)
where $|\alpha|^{2}+|\beta|^{2}=1$. The expectation value of $\hat{A}$ for $|\psi_{2}\rangle$ is
$\langle A\rangle=|\alpha|^{2}\underbrace{\langle n|\hat{A}|n\rangle}_{0}+|\beta|^{2}\underbrace{\langle k|\hat{A}|k\rangle}_{0}+2\operatorname{Re}(\alpha^{*}\beta)\langle n|\hat{A}|k\rangle.$ (E2)
Expanding out $\langle A\rangle$ in terms of $Q_{nk}$ yields
$\langle A\rangle=2\operatorname{Re}(\alpha^{*}\beta)Q_{nk}\left(1+e^{\frac{2\pi i}{3}(n-k)}+e^{\frac{4\pi i}{3}(n-k)}\right).$ (E3)
The sum of exponentials is 3 when $n-k=0\operatorname{mod}3$ and 0 otherwise. So, $\langle A\rangle$ is clearly maximised when (nk)(n-k) is a multiple of 3. Moreover, $|\alpha|^{2}+|\beta|^{2}=1$ puts a constraint on the maximum value of $\operatorname{Re}(\alpha^{*}\beta)$. Intuitively, the interference term will be maximised if the state is an equal superposition of $|n\rangle$ and $|k\rangle$. Therefore, the maximum value of $\operatorname{Re}(\alpha^{*}\beta)$ is $\frac{1}{2}$. To be precise, consider the following short proof: Let $\alpha=ce^{i\theta}$ and $\beta=se^{i\phi}$. Then $\operatorname{Re}(\alpha^{*}\beta)=cs\cos(\phi-\theta)$. $\cos(\phi-\theta)$ is maximised if $\phi=\theta$. Thus we can choose $\alpha=c$ and $\beta=s$, up to a global phase. Then, the constraint $c^{2}+s^{2}=1$ means we can parametrise $c$ and $s$ as $c=\cos(\xi)$ and $s=\sin(\xi)$. Therefore, $cs=\frac{1}{2}\sin(2\xi)$, which is maximised if $\sin(2\xi)=1$. Hence, $\operatorname{max}(cs)=\operatorname{max}\left(\operatorname{Re}(\alpha^{*}\beta)\right)=\frac{1}{2}$. Thus, $\langle A\rangle$ for an equal superposition of $|\psi_{2}\rangle=\frac{1}{\sqrt{2}}(|n\rangle+|k\rangle)$ with $n-k=0\operatorname{mod}3$ is
$\langle A\rangle=2\left(\frac{1}{2}\right)Q_{nk}(3)=3Q_{nk}.$ (E4)
We see that the problem of showing that $|\langle A\rangle|<1$ reduces to showing that
$|Q_{nk}|<\frac{1}{3}\quad\forall n,k\quad\text{s.t}\quad k-n=3m,\text{ for }m\in\mathbb{N}\ ,$ (E5)
where we only need to consider $m\geq 0$ because $Q_{nk}$ is symmetric. One might conjecture that the $Q_{03}$ matrix element will be the largest in magnitude that meets the kn=3mk-n=3m condition. If that holds then our proof would be complete, as $|Q_{03}|\approx 0.32<\frac{1}{3}$ (computed using Eq. C9). Therefore, the strategy we choose to take is to show two things about $Q_{nk}$:
  1. 1.
    The largest matrix element, in magnitude, on a diagonal i.e for fixed $m$, $Q_{n,n+2m+1}$ is $Q_{0,2m+1}$.
[…]
  2. 2.
    The non-zero matrix elements decrease in magnitude across the 0th row i.e $Q_{0,2m+1}$ is a decreasing function in $m$.
[…]
We already have established that the only non-zero matrix elements of the $Q_{nk}$ are those with opposite parity, and we ignore elements of the same parity when claiming these sequences are decreasing. Once those two statements are proven, we will have shown that $Q_{01}$ is the largest matrix element. This does not satisfy the condition kn=3mk-n=3m, though, and so we take the next largest matrix element, $Q_{03}$, thereby completing the proof.
E.2 Diagonals of the $Q_{nk}$ matrix
Showing the first statement is slightly complicated by the fact that the even $n$ and odd $n$ matrix elements have different formulas. When $n$ is even (and hence $k$ is odd) the relevant formula (obtained using Equations C9 and C11) is
$|Q^{\text{even}}_{nk}|=\left(\sqrt{\frac{2}{\pi}}\frac{(n-1)!!}{\sqrt{n!}}\right)\frac{(k+1)!}{(k-n)\sqrt{k!}2^{\frac{k+1}{2}}\left(\frac{k+1}{2}\right)!}.$ (E8)
When $n$ is odd (and hence $k$ is even) the relevant formula is
$|Q^{\text{odd}}_{nk}|=\left(\sqrt{\frac{2}{\pi}}\frac{n!!}{\sqrt{n!}}\right)\frac{k!}{(k-n)\sqrt{k!}2^{\frac{k}{2}}\left(\frac{k}{2}\right)!}.$ (E9)
To resolve this complication we consider the odd and even indexed subsequences separately. We define the sequence $\{a_{n}\}_{n=0}^{\infty}$, where $a_{n}=|Q_{n,n+2m+1}|$ for fixed $m$. Then, we just need to show the even indexed subsequence is decreasing ($a_{2j+2}<a_{2j}$) and that when the next element of the sequence is odd-indexed it is less than the previous even-indexed element ($a_{2j+1}<a_{2j}$).
Even Subsequence is Decreasing:
Consider the fraction $a_{2j+2}/a_{2j}$ and show it is less than 1.
$\frac{a_{2j+2}}{a_{2j}}=\left|\frac{Q_{2j+2,2j+2m+3}}{Q_{2j,2j+2m+1}}\right|$ (E10)
$|Q_{2j+2,2j+2m+3}|=\frac{(2j+1)!!}{\sqrt{(2j+2)!}}\frac{(2j+2m+4)!}{(2m+1)\sqrt{(2j+2m+3)!}2^{j+m+2}(j+m+2)!}$ (E11)
$|Q_{2j,2j+2m+1}|=\frac{(2j-1)!!}{\sqrt{(2j)!}}\frac{(2j+2m+2)!}{(2m+1)\sqrt{(2j+2m+1)!}2^{j+m+1}(j+m+1)!}$ (E12)
Therefore,
$\frac{a_{2j+2}}{a_{2j}}$ $=\frac{(2j+1)}{\sqrt{(2j+2)(2j+1)}}\frac{(2j+2m+4)(2j+2m+3)}{\sqrt{(2j+2m+3)(2j+2m+2)}2(j+m+2)}$ (E13)
$=\sqrt{\frac{(4j^{2}+4jm+8j)+2m+3}{(4j^{2}+4jm+8j)+4m+4}}$ (E14)
$\frac{a_{2j+2}}{a_{2j}}$ $<1\ .\qed$ (E15)
The Next Odd Element is less than the Preceding Even Element:
Using a very similar argument to above, consider the fraction $a_{2j+1}/a_{2j}$ and show it is less than 1.
$\frac{a_{2j+1}}{a_{2j}}=\left|\frac{Q_{2j+1,2j+2m+2}}{Q_{2j,2j+2m+1}}\right|$ (E16)
$|Q_{2j+1,2j+2m+2}|=\frac{(2j+1)!!}{\sqrt{(2j+1)!}}\frac{(2j+2m+2)!}{(2m+1)\sqrt{(2j+2m+2)!}2^{j+m+1}(j+m+1)!}$ (E17)
$|Q_{2j,2j+2m+1}|=\frac{(2j-1)!!}{\sqrt{(2j)!}}\frac{(2j+2m+2)!}{(2m+1)\sqrt{(2j+2m+1)!}2^{j+m+1}(j+m+1)!}$ (E18)
Therefore,
$\frac{a_{2j+1}}{a_{2j}}$ $=\left(\frac{(2j+1)!!}{(2j-1)!!}\right)\sqrt{\frac{(2j)!}{(2j+1)!}}\sqrt{\frac{(2j+2m+1)!}{(2j+2m+2)!}}$ (E19)
$=\sqrt{\frac{(2j+1)}{(2j+1)+2m+1}}$ (E20)
$\frac{a_{2j+1}}{a_{2j}}$ $<1\qed$ (E21)
E.3 0th Row of the $Q_{nk}$ Matrix is Decreasing
The second thing to show is that $Q_{0,2m+1}$ is a decreasing function in $m$. To accomplish this, consider the fraction $|Q_{0,2m+3}/Q_{0,2m+1}|$ and show it is less than 1.
$Q_{0,2m+3}=\left(\sqrt{\frac{2}{\pi}}\frac{(-1)!!}{\sqrt{0!}}\right)\frac{(2m+4)!}{(2m+3)\sqrt{(2m+3)!}2^{m+2}\left(m+2\right)!}$ (E22)
$Q_{0,2m+1}=\left(\sqrt{\frac{2}{\pi}}\frac{(-1)!!}{\sqrt{0!}}\right)\frac{(2m+2)!}{(2m+1)\sqrt{(2m+1)!}2^{m+1}\left(m+1\right)!}$ (E23)
where (1)!!=1(-1)!!=1. Therefore,
$\left|\frac{Q_{0,2m+3}}{Q_{0,2m+1}}\right|$ $=\frac{2m+1}{\sqrt{(2m+3)(2m+2)}}<\frac{2m+1}{\sqrt{(2m+1)(2m+1)}}$ (E24)
$\left|\frac{Q_{0,2m+3}}{Q_{0,2m+1}}\right|$ $<1\qed$ (E25)
Appendix F Calculation of correlators in the quantum harmonic oscillator
Consider a general state
$|\psi\rangle=\sum_{n=0}^{N}c_{n}|n\rangle,\ c_{n}\in\mathbb{C}\ .$ (F1)
The measurement operator in the Heisenberg picture is given by $\hat{Q}(t)=e^{iHt}(2\hat{\theta}-1)e^{-iHt}$, where we choose $t_{2}-t_{1}=\tau$ and $t_{1}=0$, such that
$\hat{Q}_{j}=e^{iH(j-1)\tau}(2\hat{\theta}-1)e^{-iH(j-1)\tau}\ .$ (F2)
Then,
$C_{12}=\frac{1}{2}\langle\psi|\{\hat{Q}_{1},\hat{Q}_{2}\}|\psi\rangle\ .$ (F3)
$\therefore C_{12}=\frac{1}{2}\sum_{n,k=0}^{N}c_{n}^{*}c_{k}\langle n|\hat{Q}_{1}e^{i\hat{H}\tau}\hat{Q}_{1}e^{-i\hat{H}\tau}|k\rangle+\frac{1}{2}\sum_{n,k=0}^{N}c_{n}^{*}c_{k}\langle n|e^{i\hat{H}\tau}\hat{Q}_{1}e^{-i\hat{H}\tau}\hat{Q}_{1}|k\rangle$ (F4)
We can write
$e^{i\hat{H}\tau}=\sum_{m=0}^{\infty}e^{iE_{m}\tau}|m\rangle\langle m|\ ,$ (F5)
so that
$C_{12}=\frac{1}{2}\sum_{n,k=0}^{N}c_{n}^{*}c_{k}\langle n|\hat{Q}_{1}\sum_{m=0}^{\infty}e^{iE_{m}\tau}|m\rangle\langle m|\hat{Q}_{1}e^{-iE_{k}\tau}|k\rangle+\frac{1}{2}\sum_{n,k=0}^{N}c_{n}^{*}c_{k}\langle n|e^{iE_{n}\tau}\hat{Q}_{1}\sum_{m=0}^{\infty}e^{-iE_{m}\tau}|m\rangle\langle m|\hat{Q}_{1}|k\rangle\ ,$ (F6)
which can be reduced to a form containing only the $Q_{nk}\equiv\langle n|\hat{Q}_{1}|k\rangle$ matrix elements:
$\frac{1}{2}\sum_{n,k=0}^{N}\sum_{m=0}^{\infty}e^{i(E_{m}-E_{k})\tau}c_{n}^{*}c_{k}Q_{nm}Q_{mk}+\frac{1}{2}\sum_{n,k=0}^{N}\sum_{m=0}^{\infty}e^{i(E_{n}-E_{m})\tau}c_{n}^{*}c_{k}Q_{nm}Q_{mk}\ .$ (F7)
The matrix $\mathbb{J}$ is real and symmetric; and so we can combine terms as:
$\frac{1}{2}\sum_{n,k=0}^{N}\sum_{m=0}^{\infty}e^{i(E_{m}-E_{k})\tau}c_{n}^{*}c_{k}Q_{nm}Q_{mk}+\frac{1}{2}[\sum_{n,k=0}^{N}\sum_{m=0}^{\infty}e^{i(E_{m}-E_{k})\tau}c_{n}^{*}c_{k}Q_{nm}Q_{mk}]^{*}\ .$ (F8)
That is,
$\mathrm{Re}\{\sum_{n,k=0}^{N}\sum_{m=0}^{\infty}e^{i(E_{m}-E_{k})\tau}c_{n}^{*}c_{k}Q_{nm}Q_{mk}\}\ .$ (F9)
Taking $E_{n}=n\omega$ (with $\hbar=1$), ignoring the vacuum contribution $\frac{1}{2}\hbar\omega$, we obtain
$C_{12}=\mathrm{Re}\{\sum_{n,k=0}^{N}\sum_{m=0}^{\infty}e^{i(m-k)\omega\tau}c_{n}^{*}c_{k}Q_{nm}Q_{mk}\}\ .$ (F10)
In fact, for states like the Zaw-Scarani state with support entirely on a given $\mathcal{H}_{k}$, it is sufficient to just compute $C_{12}$ as, with a measurement time spacing of a third of a period, $C_{12}=C_{13}=C_{23}=C$. We prove this statement in Appendix G. For more general states, the calculation procedure for $C_{23}$ and $C_{13}$ is very similar to Eq.F10.
Appendix G All Three Correlators are Equal on $\mathcal{H}_{k}$
For some state $|\psi_{k}\rangle\in\mathcal{H}_{k}$, the correlators $C_{ij}$ can be written as
$C_{ij}=\frac{1}{2}\langle\psi_{k}|\{\hat{Q}_{i},\hat{Q}_{j}\}|\psi_{k}\rangle\ .$ (G1)
Now, the expectation value $F_{ijk}=\frac{1}{2}\langle\psi_{k}|\{\hat{Q}_{i},\hat{Q}_{j}\}|\psi_{k}\rangle$, $i<j$, can be expanded out as
[…]
by leveraging the unitarity of $\hat{U}$ and its action on $|\psi_{k}\rangle$. We then group some terms in the aim of showing $F_{12k}=F_{23k}=F_{13k}$,
$F_{ijk}=\operatorname{Re}\omega^{k(j-i)}\langle\psi_{k}|\hat{Q}_{1}\hat{U}^{i-j}\hat{Q}_{1}|\psi_{k}\rangle\ .$ (G3)
Clearly, $F_{12k}=\operatorname{Re}\omega^{k}\langle\psi_{k}|\hat{Q}_{1}\hat{U}^{-1}\hat{Q}_{1}|\psi_{k}\rangle=F_{23k}$ but it also quickly follows that $F_{12k}=F_{13k}$. Note,
$F_{13k}=\operatorname{Re}\omega^{2k}\langle\psi_{k}|\hat{Q}_{1}\hat{U}^{-2}\hat{Q}_{1}|\psi_{k}\rangle\ ,$ (G4)
where $\omega^{2k}=(\omega^{k})^{*}$, $\hat{U}^{-2}=\hat{U}$ and $\langle\psi_{k}|\hat{Q}_{1}\hat{U}^{-2}\hat{Q}_{1}|\psi_{k}\rangle=\langle\psi_{k}|\hat{Q}_{1}\hat{U}^{-1}\hat{Q}_{1}|\psi_{k}\rangle^{*}$. So,
$F_{13k}=\operatorname{Re}\left(\omega^{k}\langle\psi_{k}|\hat{Q}_{1}\hat{U}^{-1}\hat{Q}_{1}|\psi_{k}\rangle\right)^{*}=\operatorname{Re}\omega^{k}\langle\psi_{k}|\hat{Q}_{1}\hat{U}^{-1}\hat{Q}_{1}|\psi_{k}\rangle=F_{12k}=F_{23k}\ .$ (G5)
Hence, by pairing Eq. G1 and Eq. G5, we have shown that for a given quantum state, with support entirely on $\mathcal{H}_{k}$, all the correlators are equal when the measurement spacing is $\Delta t=2\pi/3$.
Appendix H Calculation of sequential measurement probabilities
We consider the initial density operator
$\rho=|\psi\rangle\langle\psi|=\sum_{n,k=0}^{N}c_{n}^{*}c_{k}|k\rangle\langle n|$ (H1)
and define the Heisenberg-picture projection operators $P_{+}(t)=\hat{\theta}(t)$ and $P_{-}(t)=(1-\hat{\theta}(t))$, where we have defined the time-evolved Heaviside projector onto the positive x-axis:
$\hat{\theta}(t)\equiv\int_{0}^{\infty}e^{-iHt}|x\rangle\langle x|e^{iHt}\mathrm{d}x\ .$ (H2)
Then we have the simple expression
$p_{123}(s_{1},s_{2},s_{3})=\mathrm{Tr}\left[P_{s_{3}}(t_{3})P_{s_{2}}(t_{2})P_{s_{1}}(t_{1})\rho P_{s_{1}}(t_{1})P_{s_{2}}(t_{2})\right]\ .$ (H3)
In order to compute these probabilities numerically, however, it is easiest to dissociate the time dependence of the operators such that we recover expressions in terms of the familiar $Q_{nk}$ matrix elements. That is,
$p_{123}(+,+,+)=\mathrm{Tr}(e^{2iH\tau}P_{+}(t_{1})e^{-iH\tau}P_{+}(t_{1})e^{-iH\tau}P_{+}(t_{1})\rho P_{+}(t_{1})e^{iH\tau}P_{+}(t_{1})e^{-iH\tau})\ .$ (H4)
Then, writing
$e^{iH\tau}=\sum_{m=0}^{\infty}e^{im\omega\tau}|m\rangle\langle m|\ ,$ (H5)
setting $\omega=1$ as always, we define the matrix elements
$L_{nk}\equiv\langle n|(\hat{Q}_{1}+1)|k\rangle=Q_{nk}+\delta_{nk}\ ,$ (H6)
where $\delta_{nk}$ is the Kronecker delta symbol, such that
$p_{123}(+,+,+)=\frac{1}{32}\sum_{m,p,q,r=0}^{\infty}\sum_{n,k=0}^{N}c_{n}^{*}c_{k}e^{i\tau(q+r-m-p)}L_{rm}L_{mp}L_{pk}L_{nq}L_{qr}\ ,$ (H7)
recalling that $P_{+}(t_{1})=\frac{(\hat{Q}_{1}+1)}{2}$. Similarly,
$p_{123}(-,-,-)=\frac{1}{32}\sum_{m,p,q,r=0}^{\infty}\sum_{n,k=0}^{N}c_{n}^{*}c_{k}e^{i\tau(q+r-m-p)}L^{\prime}_{rm}L^{\prime}_{mp}L^{\prime}_{pk}L^{\prime}_{nq}L^{\prime}_{qr}\ ,$ (H8)
where $\mathbf{L}^{\prime}\equiv(\mathbf{\mathbb{1}}-\mathbf{Q})$. Computationally, we have chosen a finite cut-off value $N$ to truncate the infinite sum over $m,p,q,r$.
We can alternatively consider making the first measurement a “weak measurement”, which leads to the expression
$p_{(1)23}(+,+,+)=\frac{1}{2}\text{Tr}\left(P_{+}(t_{3})P_{+}(t_{2})(P_{+}(t_{1})\rho+\rho P_{+}(t_{1}))P_{+}(t_{2})\right)=\frac{1}{2}\text{Tr}\left(P_{+}(t_{3})P_{+}(t_{2})\left\{P_{+}(t_{1}),\rho\right\}P_{+}(t_{2})\right)\ .$ (H9)
Classically, this distribution is identical to $p_{123}(+,+,+)$; but quantum mechanically, the sum rule (which is not generally valid) does indeed hold over the weak first measurement:
$\sum_{s_{1}}p_{23}(s_{1},s_{2},s_{3})=p_{23}(s_{2},s_{3})\ .$ (H10)
This distribution therefore could be said to lie between the intrusive three-time measurement probability, $p_{123}$, and the quasi-probability distribution, $q$ (where the latter is equivalent to making all three measurements weak). In this case,
$p_{(1)23}(+,+,+)-p_{(1)23}(-,-,-)\equiv\Delta p_{(1)23}=\frac{1}{4}(\langle A\rangle+D+I)\ ,$ (H11)
where the interference terms would only stem from the second measurement: $I=\langle Q_{3}^{(2)}\rangle-\langle Q_{3}\rangle$. We could just as well consider a distribution with a weak second measurement, where
$p_{1(2)3}(+,+,+)=\frac{1}{2}\text{Tr}\left(P^{(3)}_{+}\{P^{(2)}_{+},P^{(1)}_{+}\rho P^{(1)}_{+}\}\right)\ ,$ (H12)
and so
$p_{1(2)3}(+,+,+)=\frac{1}{2}\text{Tr}\left(e^{2iH\tau}\hat{\theta}e^{-2iH\tau}\{e^{iH\tau}\hat{\theta}e^{-iH\tau},\hat{\theta}\rho\hat{\theta}\}\right)\\ =\frac{1}{2}\text{Tr}\left(e^{2iH\tau}\hat{\theta}e^{-iH\tau}\hat{\theta}e^{iH\tau}\hat{\theta}\rho\hat{\theta}\right)+\frac{1}{2}\text{Tr}\left(e^{2iH\tau}\hat{\theta}e^{-2iH\tau}\hat{\theta}\rho\hat{\theta}e^{iH\tau}\hat{\theta}e^{-iH\tau}\right)\ ,$ (H13)
which can be computed straightforwardly in terms of the afore-defined $L_{nk}$ matrix elements; for example,
$\frac{1}{2}\text{Tr}\left(e^{2iH\tau}\hat{\theta}e^{-iH\tau}\hat{\theta}e^{iH\tau}\hat{\theta}\rho\hat{\theta}\right)=\frac{1}{32}\sum_{n,k,p,q,r}c_{n}^{*}c_{k}e^{i\tau(2p+r-q)}L_{pq}L_{qm}L_{mk}L_{np}\ .$ (H14)
Appendix I Tsirelson Violation implies LG3 satisfaction on $\mathcal{H}_{k}$
I.1 Outline and Preliminaries
This appendix is devoted to proving that if Tsirelson’s inequality is violated by a given quantum state, with support entirely in $\mathcal{H}_{k}$, then that state will satisfy all four LG3s (Eqs 1.12-1.15) if the measurement times $t_{1},t_{2},t_{3}$ are the same as in the Tsirelson scheme. This result highlights how the Tsirelson test probes nonclassicality in a different way than the standard LG3s and begins to connect the two frameworks.
One simply needs to show two statements hold true. The conjunction of these two statements would imply that all four LG3s are greater than zero:
  1. 1.
    A Tsirelson violation implies satisfaction of the first LG3 i.e $L_{1}>0$. We straightforwardly show this in Section I.2. In fact, this result holds for arbitrary quantum states.
  2. 2.
    For quantum states in $\mathcal{H}_{0},\mathcal{H}_{1},\mathcal{H}_{2}$, the first Leggett–Garg quantity is always the smallest of the four. Consequently, all four LG3s are positive.
I.2 Tsirelson Violation $\implies L_{1}>0$
Consider a state that violates Tsirelson’s inequality such that
$\langle Q_{1}+Q_{2}+Q_{3}\rangle>1$ (I1)
or in fact $\langle Q_{1}+Q_{2}+Q_{3}\rangle<-1$ (this doesn’t change the argument at all). Now, consider the Cauchy-Schwarz inequality applied to the left hand side of Equation I1
$\langle(Q_{1}+Q_{2}+Q_{3})^{2}\rangle\geq(\langle Q_{1}+Q_{2}+Q_{3}\rangle)^{2}.$ (I2)
Using Equations I1 and I2 we can write
$\langle(Q_{1}+Q_{2}+Q_{3})^{2}\rangle>1.$ (I3)
Finally, simply expanding out the expectation value nets
$1+C_{12}+C_{13}+C_{23}>0.$ (I4)
I.3 $L_{1}$ is the smallest LG3 on $\mathcal{H}_{0},\mathcal{H}_{1}$ and $\mathcal{H}_{2}$
Recall from Appendix G that for $|\psi_{k}\rangle\in\mathcal{H}_{k}$, $C_{12}=C_{13}=C_{23}=C$. Therefore, on these subspaces, $L_{1}=1+3C$ and the rest are 12C1-2C. So, by showing that $C_{12}$ is negative we will have shown that $L_{1}$ is the smallest LG3. Note,
$C_{12}=\frac{1}{2}\langle\psi_{k}|Q(\omega^{-k}U+\omega^{k}U^{\dagger})Q|\psi_{k}\rangle\ ,$ (I5)
where $Q|\psi_{k}\rangle$ can be decomposed into $Q|\psi_{k}\rangle=|\phi_{k}\rangle+|\phi_{\perp}\rangle$, $|\phi_{k}\rangle\in\mathcal{H}_{k}$ and $\langle\phi_{k}|\phi_{\perp}\rangle=0$. Hence, as $|\phi_{\perp}\rangle$ is constructed out of orthogonal energy eigenstates to $|\phi_{k}\rangle$, we can write
$C_{12}=\frac{1}{2}\langle\phi_{k}|(\omega^{-k}U+\omega^{k}U^{\dagger})|\phi_{k}\rangle+\frac{1}{2}\langle\phi_{\perp}|(\omega^{-k}U+\omega^{k}U^{\dagger})|\phi_{\perp}\rangle\ .$ (I6)
Motivated by the roots of unity representation of the cyclic group, note that
$(\mathds{1}+\omega^{-k}U+\omega^{k}U^{\dagger})|\phi_{k}\rangle=3|\phi_{k}\rangle\Rightarrow(\omega^{-k}U+\omega^{k}U^{\dagger})|\phi_{k}\rangle=2|\phi_{k}\rangle\ ,$ (I7)
and
$(\mathds{1}+\omega^{-k}U+\omega^{k}U^{\dagger})|\phi_{\perp}\rangle=0\Rightarrow(\omega^{-k}U+\omega^{k}U^{\dagger})|\phi_{\perp}\rangle=-|\phi_{\perp}\rangle\ .$ (I8)
Thus, we can write Equation I6 as
$C_{12}=\langle\phi_{k}|\phi_{k}\rangle-\frac{1}{2}\langle\phi_{\perp}|\phi_{\perp}\rangle\ .$ (I9)
By substituting in the condition $\langle\phi_{\bot}|\phi_{\bot}\rangle+\langle\phi_{k}|\phi_{k}\rangle=1$ we find
$C_{12}=\frac{3}{2}\langle\phi_{k}|\phi_{k}\rangle-\frac{1}{2}\ .$ (I10)
Reintroducing $|\psi_{k}\rangle$ nets an inequality including the spectral norm of an operator
[…]
where $P_{k}$ is the projection operator onto $\mathcal{H}_{k}$. We can trivially extend this upper bound by writing
[…]
But,
[…]
Therefore,
[…]
Finally, using the upper bound on $A||$ from Ref. [9],
$C_{12}\leq\frac{1}{6}(1.39)^{2}-\frac{1}{2}\leq 0\ .$ (I15)
Appendix J Constrained Optimisation Procedure for Method 2
We seek a state that violates the Tsirelson inequality but satisfies the condition $p_{123}(+,+,+)=p_{123}(-,-,-)$, or $\Delta p_{123}=0$, exactly. We can reframe this question as a constrained optimisation problem where the function we wish to maximise is
$\langle A\rangle\equiv\langle\hat{Q}_{1}+\hat{Q}_{2}+\hat{Q}_{3}\rangle$ (J1)
which can be written as a function of the coefficients in the energy basis expansion of the initial state:
$\langle A\rangle=f(c_{n}).$ (J2)
Recall that
$p_{123}(+,+,+)=\frac{1}{32}\sum_{m,p,q,r=0}^{\infty}\sum_{n,k=0}^{N}c_{n}^{*}c_{k}e^{i\tau(q+r-m-p)}L_{rm}L_{mp}L_{pk}L_{nq}L_{qr}\ ,$ (J3)
where the $L_{nk}$ matrix elements are defined as in Eq. H6, with a similar expression for $p_{123}(-,-,-)$. Defining
$\Gamma_{nk}=L_{rm}L_{mp}L_{pk}L_{nq}L_{qr}\ ,$ (J4)
the constraint can be rewritten in the simple form
$\Delta p_{123}\equiv 2\sum_{n,k\text{ with opposite parity}}^{N}c_{n}^{*}c_{k}\Gamma_{nk}=0\ .$ (J5)
or more algorithmically:
$\sum_{j=1}^{\left\lfloor\frac{N+1}{2}\right\rfloor}\sum_{i=0}^{N-(2j-1)}\text{Re}(c_{i}^{*}c_{i+2j-1}\Gamma_{i,i+2j-1})=0\ .$ (J6)
Thus, the objective function that we will optimise is
$L(c_{n},\lambda)=f(c_{n})-\lambda\left(\sum_{j=1}^{\left\lfloor\frac{N+1}{2}\right\rfloor}\sum_{i=0}^{N-(2j-1)}\text{Re}(c_{i}^{*}c_{i+2j-1}\Gamma_{i,i+2j-1})\right),$ (J7)
where the constraint is introduced as a Lagrange multiplier. We also introduce a constraint on the norm of the vector of being 1.
We find that the state
$(-0.281+0.52i)|0\rangle+(0.096-0.127i)|1\rangle+(-0.024+0.102i)|2\rangle\\ +(0.423-0.539i)|3\rangle+(-0.066+0.068i)|4\rangle+(0.012-0.067i)|5\rangle\\ +(-0.25+0.262i)|6\rangle$ (J8)
readily produces a Tsirelson violation of $\langle A\rangle=1.0756$ with $\Delta p_{123}=0$. More accurately, $\Delta p_{123}$ is not precisely and analytically zero for the quoted state, since the numerical computation was done using a truncated energy expansion. However, we find that slightly tweaking the state parameters can lead to states with $\Delta p_{123}$ slightly above or below zero, and so it is almost certain that, in principle, there exists a state with $\Delta p_{123}=0$ exactly.
Appendix K Probability Currents
K.1 Relating the Tsirelson quantity to the quantum current
Consider the limit definition of the integral to write $\langle\hat{T}_{D}\rangle$ as a Riemann sum,
$\langle\hat{T}_{D}\rangle$ $=\frac{1}{2\pi}\left[\lim_{N\rightarrow\infty}\frac{2\pi}{N+1}\sum_{n=0}^{N}\left\langle\theta\left(\frac{2\pi n}{N+1}\right)\right\rangle\right]\ ,$ (K1)
$=\frac{1}{3}\sum_{n=0}^{2}\left\langle\theta\left(\frac{2\pi n}{N+1}\right)\right\rangle+\frac{\varepsilon}{2\pi}\ ,$ (K2)
where $\varepsilon$ refers to an error term. The equality certainly holds for some $\varepsilon\in\mathbb{R}$.
Assuming $\langle\theta(t)\rangle$ is analytic in $t$, then $\langle\theta(t)\rangle$ has a well-defined Taylor expansion and we can borrow some well-known results from numerical analysis to quantify our error term. Specifically, we notice that this amounts to a left endpoint approximation to the integral for the dwell time. Therefore we can use the formula for the error term [75],
$E_{L}=\frac{(2\pi)^{2}}{6}\frac{d\langle\theta\rangle}{dt}\big{|}_{t=\tau},\tau\in[0,2\pi]\ ,$ (K3)
to state that
$\langle\hat{T}_{D}\rangle$ $=\frac{1}{3}\sum_{n=0}^{2}\left\langle\theta\left(\frac{2\pi n}{N+1}\right)\right\rangle+\frac{E_{L}}{2\pi}$ (K4)
$=\frac{1}{3}\sum_{n=0}^{2}\left\langle\theta\left(\frac{2\pi n}{N+1}\right)\right\rangle+\frac{2\pi}{6}\frac{d\langle\theta\rangle}{dt}\big{|}_{t=\tau}\ ,$ (K5)
where $\tau\in[0,2\pi]$. The sum can be related to the Tsirelson quantity because $\left\langle\theta\right\rangle=\frac{\left\langle Q\right\rangle+1}{2}$ and $\langle A\rangle=\sum_{n=0}^{2}\left\langle Q\right\rangle$. So,
$\langle\hat{T}_{D}\rangle$ $=\frac{1}{6}\left(\sum_{n=0}^{2}\left\langle Q\left(\frac{2\pi n}{N+1}\right)\right\rangle+3\right)+\frac{2\pi}{6}\frac{d\langle\theta\rangle}{dt}\big{|}_{t=\tau}$ (K6)
$=\frac{1}{6}\left(3+\langle A\rangle\right)+\frac{2\pi}{6}\frac{d\langle\theta\rangle}{dt}\big{|}_{t=\tau}.$ (K7)
Hence,
$\langle A\rangle=(6\langle\hat{T}_{D}\rangle-3)-2\pi\frac{d\langle\theta\rangle}{dt}\big{|}_{t=\tau}=-2\pi\frac{d\langle\theta\rangle}{dt}\big{|}_{t=\tau}$ (K8)
This expression thereby provide an alternative explanation for why $\langle A\rangle=0$ for energy eigenstates (stationary states). The expectation value of the Heaviside operator has no time dependence for energy eigenstates, so those derivatives are 0 for any value of $\tau$.
Note that because $\langle\theta(t+\pi)\rangle=1-\langle\theta(t)\rangle$, the derivatives of $\langle\theta(t)\rangle$ simply flip sign after half a period. Hence, without loss of generality, we may drop the negative signs on our expressions, since $\tau$ is some unknown time on the interval $0,2\pi]$:
$\langle A\rangle=2\pi\frac{d\langle\theta\rangle}{dt}\big{|}_{t=\tau}$}\ . (K9)
We can rewrite Eq. K9 as,
$\langle A\rangle=2\pi\frac{d}{dt}\int_{0}^{\infty}|\psi(x,\tau)|^{2}dx.$ (K10)
Then, using the 1D continuity equation,
$\frac{\partial\rho}{\partial t}+\frac{\partial J}{\partial x}=0,$ (K11)
where $\rho=|\psi|^{2}$ is the probability density and $J=\operatorname{Im}\left(\psi^{*}\partial_{x}\psi\right)$ is the probability density current [76], we can substitute the current into the Tsirelson quantity:
$\langle A\rangle$ $=-2\pi\int_{0}^{\infty}\frac{\partial J(x,\tau)}{\partial x}dx,$ (K12)
$=2\pi J(0,\tau).$ (K13)
Therefore, assuming $J(\infty,\tau)\rightarrow 0$, we have directly related the Tsirelson quantity to the probability current at the origin. Specifically,
$\exists\tau\in[0,2\pi]\text{ }\operatorname{s.t}\text{ }\langle A\rangle=2\pi J(0,\tau).$ (K14)
From Equation K14, we can make the following claim: if $|J(0,\tau)|\leq\frac{1}{2\pi},\forall\tau\in[0,2\pi]$, then $|\langle A\rangle|\leq 1$. That is to say, if the current at the origin stays sufficiently low, then we can guarantee the state will not violate the Tsirelson inequality. Notably, $J(0,t)$ could be measured by taking the time derivative of measured values of $\langle\theta(t)\rangle$ which only requires single-time measurements.
This result also provides some further insight into why completely even or odd wavefunctions cannot violate the Tsirelson inequality. Expanding a general state as a superposition of energy eigenstate wavefunctions $\psi_{n}(x)$, the current at the origin can be written as
$J(0,t)=\operatorname{Im}\sum_{n,m}c_{n}^{*}c_{m}\psi_{n}(0)\psi_{m}^{\prime}(0)e^{i(n-m)t}\ .$ (K15)
Even functions have 0 slope at $x=0$ and odd functions are 0 at $x=0$. As such, if the wavefunction in question can be decomposed into exclusively even energy eigenfunctions or exclusively odd energy eigenfunctions, then $J(0,t)=0,\forall t\in[0,2\pi]$, so $\langle A\rangle=0$.
K.2 A Tsirelson satisfaction condition on $\mathcal{H}_{0}$
We will refer to the dwell time over $2/3$ of a period as $T_{2/3}$. Consider the following integral approximation:
$T_{2/3}$ $=\frac{1}{6}\left(\sum_{j=1}^{3}\langle\theta_{j}\rangle+\langle\theta_{1}\rangle\right)+\frac{1}{2\pi}\sum_{j=1}^{3}E_{j}$ (K16)
$=\frac{1}{12}\langle A\rangle+\frac{1}{12}\langle Q_{1}\rangle+\frac{1}{3}+\frac{1}{2\pi}\sum_{j=1}^{3}E_{j}.$ (K17)
Therefore,
$\langle A\rangle=12T_{2/3}-4-\langle Q_{1}\rangle-\frac{6}{\pi}\sum_{j=1}^{3}E_{j}$ (K18)
where both $E_{1}$ and $E_{2}$ are left end point errors with $E_{1}=\frac{(2\pi/3)^{2}}{2}\frac{\mathrm{d}\langle\theta\rangle}{\mathrm{d}t}\big{|}_{\tau_{1}}\ ,\ \tau_{1}\in[0,2\pi/3]$ and $E_{2}=\frac{(\pi/3)^{2}}{2}\frac{\mathrm{d}\langle\theta\rangle}{\mathrm{d}t}\big{|}_{\tau_{2}}\ ,\ \tau_{2}\in[2\pi/3,\pi]$. However, $E_{3}$ is a right endpoint error with $E_{3}=-\frac{(\pi/3)^{2}}{2}\frac{\mathrm{d}\langle\theta\rangle}{\mathrm{d}t}\big{|}_{\tau_{3}}\ ,\ \tau_{3}\in[\pi,4\pi/3]$. So, expanding out and simplifying the error term contribution,
$\sum_{j=1}^{3}E_{j}=\frac{\pi^{2}}{18}\left[4J(\tau_{1})+J(\tau_{2})-J(\tau_{3})\right]\ .$ (K19)
Therefore,
$\langle A\rangle=12T_{2/3}-4-\langle Q_{1}\rangle-\frac{\pi}{3}\left[4J(\tau_{1})+J(\tau_{2})-J(\tau_{3})\right]\ .$ (K20)
But, by noting that $J(\tau_{3})=+J(\tilde{\tau}_{3}),\ \tilde{\tau}_{3}\in[0,\pi/3]$, the sum of the three currents can be reformulated as a weighted average over the domain $0,\pi]$. Pairing that with an application of the mean value theorem allows us to write
$\langle A\rangle=12T_{2/3}-4-\langle Q_{1}\rangle-2\pi J(\tau),\ \tau\in[0,\pi]\ .$ (K21)
For $|\psi\rangle\in\mathcal{H}_{0}$, we have that $T_{2/3}=1/3$ and $\langle Q_{1}\rangle=\frac{1}{3}\langle A\rangle$. Therefore, the expression for $\langle A\rangle$ simplifies to
$\langle A\rangle=-\frac{3\pi}{2}J(\tau),\ \tau\in[0,2\pi/3]\ ,$ (K22)
where the domain of $\tau$ was decreased as the current is periodic over $2\pi/3$. Once again an application of the sign flipping property of the current allows us to say
$\text{For}\ |\psi\rangle\in\mathcal{H}_{0}\ ,\ \exists\tau\in[0,2\pi/3]\ \text{s.t}\ \langle A\rangle=\frac{3\pi}{2}J(\tau).$ (K23)
An immediate corollary of this is that
$\text{For}\ |\psi\rangle\in\mathcal{H}_{0}\ ,\ \text{if}\ |J(\tau)|\leq\frac{2}{3\pi}\ \forall\tau\in[0,2\pi/3]\ \text{then}\ |\langle A\rangle|\leq 1\ .$ (K24)
Curiously, when we computed the classical probability current, using
$j_{c}(0,t)=\iint\mathrm{d}x\mathrm{d}p\ p(t)\delta[x(t)]W_{c}(x,p)\ ,$ (K25)
for any natural classicalisation of the phase-space distribution, $W_{c}(x,p)$, associated with $|\psi_{Z}\rangle\in\mathcal{H}_{0}$, we found the classical current was lower in magnitude than the quantum current across the entire period. Specifically, from the Wigner function of the Zaw-Scarani state, $W_{Z}(x,p)$, we constructed the following classical distributions: (1) $W_{c}(x,p)\propto|\langle\alpha|\psi_{Z}\rangle|^{2}$ (the Husimi-Q distribution), (2) $W_{c}(x,p)\propto|W_{Z}(x,p)|$, (3) $W_{c}(x,p)\propto W_{Z}(x,p)+|W_{Z}(x,p)|$ and (4) $W_{c}(x,p)=|\psi_{Z}(x)|^{2}|\phi_{Z}(p)|^{2}$ (where $\phi_{Z}(p)$ is the momentum-space wavefunction of the Zaw-Scarani state). In Figure 9, we illustrate the resulting classical current, where the classicalisation strategy chosen is using the Husimi-Q distribution, and compare it to the quantum current. This result perhaps suggests a connection between the current based bound on the Tsirelson quantity in Eq. K24 and classicalised currents of violating quantum states.
Figure 9: Plot of the probability density current for the Zaw-Scarani state at the origin over a period compared to the classical probability density current from the classical time evolution of the Husimi-Q representation of the Zaw-Scarani state. Notably, the classicalised current is consistently lower than the quantum current.